### v1

# TWIST: Training-free and Label-free Short Text Clustering through Iterative Vector Updating with LLMs

Abstract
In this paper, we propose a training-free and label-free method for short text clustering that can be used on top of any existing embedder. In the context of customer-facing chatbots, companies are dealing with large amounts of user utterances that need to be clustered according to their intent. In these commercial settings, no labeled data is typically available, and the number of clusters is not known. Our method is based on iterative vector updating: it constructs sparse vectors based on representative texts, and then iteratively refines them through LLM guidance. Our method achieves comparable or superior results to state-of-the-art methods that use contrastive learning, but without assuming prior knowledge of clusters or labels. Experiments on diverse datasets and smaller LLMs show that our method is model agnostic and can be applied to any embedder, with relatively small LLMs, and different clustering methods. We also show that our method scales to large datasets, reducing the computational cost of the LLM. These low-resource, adaptable settings and the scalability of our method make it more aligned with real-world scenarios than existing clustering methods. Our source code is available here: https://anonymous.4open.science/r/sparse_vector-F4C9/README.md
1 Introduction
Short text clustering is an NLP task that automatically groups unlabeled data into clusters and is widely used in practical applications for text classification and new category detection. For example, in conversational understanding, user intents in real-world scenarios are dynamic and continuously evolving (Liang et al., 2024c), and relying on expert labeling can be costly. The short-text clustering task enables companies to efficiently organize tens of thousands of client requests in real time. Other downstream tasks, such as mining topics and opinions from online platforms, also rely on effective clustering (Wu et al., 2024).
The clustering task is often approached as a Generalized Category Discovery (GCD) scenario (Vaze et al., 2022), where limited labeled data and substantial unlabeled data are used to cluster all unlabeled samples, which include both known and unknown classes. Recently, other studies have explored an emerging data scenario, where no labels are provided at all (De Raedt et al., 2023; Viswanathan et al., 2024b; Zhang et al., 2023; Lin et al., 2025). This scenario reduces reliance on human annotation and shows greater flexibility.
A common approach for the emerging data scenarios is to train an embedder with contrastive learning to improve text representations (Zhang et al., 2021). With the advancement of LLMs, training-free approaches have also been proposed (De Raedt et al., 2023). The emerging data scenario has attracted increasing attention in recent research (Zhang et al., 2023; Feng et al., 2024; Lin et al., 2025). Despite the progress of prior work, these label-free approaches still face certain limitations and leave room for further improvement. First, methods involving LLM feedback often assume known number of clusters during method development (Zhang et al., 2023; Feng et al., 2024), which is difficult to estimate in real-world settings. Second, training-free methods have mainly been explored with large, proprietary models (Viswanathan et al., 2024b; De Raedt et al., 2023), which can be costly. Third, the scalability of these methods remains a concern. Motivated by these challenges, we propose a training-free approach that requires only a smaller LLM, alleviates the known-cluster assumption, and is scalable. The core idea of our method is to construct new representations (referred to as sparse vectors in this paper) such that texts belonging to the same cluster become more similar. Our method consists of two stages: we first construct sparse representations based on representative texts, and then iteratively refine them by combining LLM guidance with pre-trained embeddings.
The contributions of our method are as follows:
  • Inclusive and easy to optimize: Builds on existing embedders and requires only some hyperparameter tuning.
  • Flexible: Does not assume a known number of clusters, nor require labeled data, contrastive learning on the task specific datasets, or fine-tuning, enabling easier adaptation to new domains.
  • Low-resource: Even with small LLMs and embedders, the results are already competitive with methods that rely on labeled data or fine-tuning.
  • Scalable: Does not require LLM feedback for every text, unlike many previous studies.
2 Related Work
Short text clustering
Generalized Category Discovery (GCD) is often studied for this task (Liang et al., 2024b; Liang and Liao, 2023; Zou et al., 2025b). These works use limited labeled text and substantial unlabeled text to learn clustering representations. However, these studies use train–test splits within a single dataset, which implicitly assumes that new categories are similar to the known ones. In practice, emerging data may show topic drift. Recently, with the advancement of LLMs, many studies have explored the emerging data scenario (Zhang et al., 2023; Viswanathan et al., 2024b; Lin et al., 2025), where no labels are available. We follow this unsupervised scenario.
Boosting Unsupervised Short Text Clustering with LLM-Guided Contrastive Learning
Contrastive learning is commonly used in unsupervised short text clustering (Zhang et al., 2021, 2023; Liang et al., 2024a). Zhang et al. (2023) use LLMs to construct hard triplets to derive training pairs, which are then used to fine-tune a small embedder. Liang et al. (2024a) use LLMs to refine texts by generating descriptors, then using small language models for contrastive learning to align the embeddings of texts and their descriptors. However, these methods either rely on knowing the number of clusters, proprietary models, or fine-tuning. Our method does not have these requirements.
Training-Free Methods for Unsupervised Short Text Clustering
Training-free approaches use a frozen pre-trained embedder: De Raedt et al. (2023) select prototypical utterances and then use an LLM to generate a set of descriptive utterance labels for these prototypes. The labels and texts are then encoded to form better representations. Viswanathan et al. (2024a) use LLMs to improve texts by generating key phrases for each text, which are then separately encoded into embeddings and concatenated for clustering. Lin et al. (2025) use LLMs for selection rather than text enrichment. They first encode all of the texts, and then use the LLM to select similar pairs for each text. Each text embedding is then obtained by average pooling the embeddings of the text and its similar pairs. These methods have in common that they refine original embeddings to improve clustering. In contrast, we construct a new vector using LLM feedback. Feng et al. (2024) use LLMs to refine edge points. However, the assumption of a known number of clusters can still be a core problem in real-world applications.
3 Theoretic formulation and assumptions
3.1 Problem definition
In our work, we consider two variations of the short text clustering task: (i) the emerging data scenario and (ii) the GCD scenario. In the emerging data scenario (Lin et al., 2025), the only available resource is a collection of unlabeled texts $D^{test}\triangleq\{x_{i}\}_{i=1}^{N}$, where each $x_{i}\in D^{test}$ corresponds to a short text and $N$ is the total number of samples. The dataset is used as a whole for clustering and evaluation, without any train-test split. In contrast, the GCD scenario has access to additional data, typically including some labeled samples (Zou et al., 2025b). In this work, we adapt this setting by considering only an additional unlabeled dataset $D^{u}\triangleq\{x_{i}\}_{i=1}^{N_{u}}$, in line with the goal of low-resource. The GCD scenario is included for the purpose of testing whether extra data improves clustering and generalizability of our method.
For both scenarios, the task is to partition $D^{test}$ into clusters $\{\hat{S}_{l}\}_{l=1}^{\hat{K}}$, where $\hat{K}$ is the number of predicted clusters. Note that $\sum_{l=1}^{\hat{K}}|\hat{S}_{l}|=N$. The goal is to ensure that the clustering results $\{\hat{S}_{l}\}_{l=1}^{\hat{K}}$ are as close as possible to the true partition $\{S_{l}\}_{l=1}^{K}$, where $K$ is the number of unique clusters in the ground truth. Importantly, unlike many existing studies (De Raedt et al., 2023; Feng et al., 2024; Zou et al., 2025b), which assume that $K$ is known in advance, we make no such assumption.
For notation convenience, we define the full dataset as $D\triangleq D^{test}\cup D^{u}$ to simplify the explanation of our proposed method. Note that $D=D^{test}$ in the emerging data scenario.
3.2 Constructing the Sparse Vector
Basic Assumption
Considering the fact that texts belonging to the same cluster are similar to each other, we represent each text as a sparse bag-of-texts vector that directly encodes its similarity to other texts. Formally, for any $x_{i},x_{j}\in D$ with $i\neq j$:
  • $x_{i}$ and $x_{j}$ are similar if $x_{i},x_{j}\in S_{k}$ for some cluster $S_{k}$.
  • $x_{i}$ and $x_{j}$ are not similar if $x_{i}\in S_{k}$ and $x_{j}\in S_{l}$ with $k\neq l$.
Construction of Bag-of-Texts
Since similarity is determined by clusters, it is not necessary to use all texts in $D$ to represent sparse vectors; instead, a representative subset of texts can constitute our bag of texts, denoted as $D^{r}\subseteq D$, where $|D^{r}|\triangleq d$. The set of representative texts $D^{r}$ should contain at least one element from each cluster. That is, for any cluster $S\in\{S_{l}\}^{K}_{l=1}$, there exists at least one text $x\in D^{r}$ such that $x\in S$.
Therefore, to cover all possible clusters, the minimum size of $d$ is K1K-1. Since $K$ is unknown, the most practical approach is to choose a relatively large $d$. For each $x\in D$, we define a $d$-diemnsional vector $\mathbf{z}$ assigned to the representative texts, where the $j$-th coordinate takes the value 1 if $x$ is similar to the $j$-th representative text, and 0 otherwise.
Proposition 1.
Given the assumption, for any $x_{i},x_{j}\in D$ with $i\neq j$ that belong to the same cluster $S_{k}$, the corresponding sparse vectors $\mathbf{z}_{i}$ and $\mathbf{z}_{j}$ will be identical.
The proposition is straightforward, so we omit the proof.
Adaptation to real-world practice
Although the proposition holds in the idealized setting, in practice similarity is typically determined using LLMs, which may introduce errors, and scanning the entire dataset for each $x$ is computationally expensive. To address these, we construct a sparse vector $\mathbf{z}$ in a way that is adaptable to real-world practice: for each $x\in D$, instead of assigning fixed value in $\{0,1\}$ based on representative texts, we propose an updating procedure that accounts for uncertainty. In addition, to address limited computational resources, we use an existing embedder to identify a relatively small set of $m$ similar candidates that are likely to belong to the same category, as empirically shown in previous research (Lin et al., 2025; Zou et al., 2025a). The details of the method will be discussed in the next section.
4 Computational Method
Our method aims to construct sparse vectors that capture the similarity of texts within the same cluster. It consists of two stages. In the initial stage, we select representative texts, which are then used to construct a sparse vector for each text. In the second stage, we iteratively update each text’s sparse vector by considering its similar texts until convergence, illustrated in Figure 1.
Unlike many prior studies that assume the number of clusters $K$ is known, we consider a more practical setting in which $K$ is unknown, and thus approximate the true number of clusters by the number of representative texts, $d$. This is less challenging than estimating $K$, as it only requires ensuring that $d$ is large enough to cover $K$ clusters.
Figure 1: Construction of Sparse vectors. Initial stage: We first encode each text in $D$ using a pretrained embedder and partition them into $d$ clusters. We then select the medoid of each cluster as the representative text and construct one-hot vectors for these representatives. The remaining text vectors are obtained based on LLM selection. Iterative stage: We update all of the vectors iteratively until convergence.
4.1 Initial stage: construction of sparse representation
To build sparse vector representations, we first need to create a set of representative samples $D^{r}$. To this end, we first encode each text using an existing embedder $F$ to obtain pre-trained embeddings:
$\mathbf{x}_{i}=F(x_{i}).$
Next, we apply a clustering algorithm to partition these embeddings into $d$ clusters.
We use agglomerative clustering (Murtagh and Legendre, 2014) as it produces deterministic results, which facilitates reproducibility. For each initial cluster , we derive the medoid. We refer to the set of medoids as the representative text, denoted as:
$D^{r}\triangleq\{x_{c_{1}},\dots,x_{c_{d}}\}.$
and let $\mathbf{X}^{r}$ be $D^{r}$ corresponding pre-trained embeddings, where $\mathbf{X}^{r}=\{\mathbf{x}_{i}:x_{i}\in D^{r}\}$. We then transform the representative texts into one-hot vectors, which serve as the basis vectors for the sparse representation.
For the remaining texts $x_{j}\in D\setminus D^{r}$, we construct their sparse vectors using this basis. For each $x_{j}$, we compute the pairwise cosine distances between its embedding $\mathbf{x}_{j}$ and the medoid embeddings in $\mathbf{X}^{r}$. By sorting these distances, we select the top $m$ closest representative texts to $x_{j}$, denoted as $x_{j1},\dots,x_{jm}$. Then, we use an LLM to identify the similar text(s) to $x_{j}$ among these representative texts, denoted as $M_{j}\triangleq LLM(x,x_{j1},\dots,x_{jm})\in\mathcal{P}(\{x_{j1},\dots,x_{jm}\})\cup\{\text{None}\}$, where $\mathcal{P}(\cdot)$ represents the power set, i.e., all possible subsets of the candidate texts, and ‘None’ indicates that no representative text is selected. If the LLM output is not None, the sparse vector of $x_{j}$ is then derived by averaging the sparse vectors of the texts in $M_{j}$, denoted as $\mathbf{z}_{j}$. Instead of assigning a fixed value of 1 to the selected dimension, we average the vectors to capture how strongly each representative text contributes to the feature. We then apply L2 normalization to ensure a consistent scale for the iterative stage. If the LLM output is None, we make this text into a representative text itself and increase the dimensionality of all sparse vectors by one and construct its corresponding one-hot vectors.
Note: if we set $d=N$, all texts will be representative texts and one-hot encoded.
4.2 Iterative stage: sparse vectors refinements
In this stage, we update the sparse vectors of all texts using the dynamic subset of $m$ candidates.
At each iteration step $t$ (with the initial stage being $t=0$), for each $x_{j}\in D$, we first concatenate its pre-trained embedding and sparse vector: $\mathbf{u}_{j}^{t}\triangleq\mathbf{x}_{j}\oplus\mathbf{z}_{j}^{t}$. We then compute the pairwise cosine distances between $\mathbf{u}_{j}^{t}$ and the concatenated representations of the all texts $\mathbf{U}^{t}$. By sorting these distances, we select the top $m$ closest texts to $x_{j}$, denoted as $x^{t}_{j1},\dots,x^{t}_{jm}$. The concatenation ensures that the pre-trained embedding guides the sparse vectors. Also, the reason we do iterative updating instead of just once is to gain more stable sparse vectors.
Next, we use an LLM to identify the subset of the texts most similar to $x_{j}$, denoted as $M^{t}_{j}\triangleq LLM(x_{j},x^{t}_{j1},\dots,x^{t}_{jm})\in\mathcal{P}(\{x^{t}_{j1},\dots,x^{t}_{jm}\})\cup\{\text{None}\}$. If the LLM output is None, we fix sparse vectors, i.e., $\mathbf{z}_{j}^{t+1}=\mathbf{z}_{j}^{t}$. Otherwise, we update the sparse vectors by averaging the sparse vector of the texts in $M^{t}_{j}$. We then apply L2 normalization, as explained previously.
For $x_{j}$, the update process is ended if the sparse vector at iteration $t+1$ is almost identical to those at iteration $t$, as we define by a cosine similarity greater than 0.99. For efficiency, we set $m=30$ and cap the maximum number of iterations at 10.
5 Experimental Settings
Emerging data $K$ $|D^{test}|$ $|D^{u}|$
Bank77 77 3,080 -
Clinc150 150 4,500 -
Mtop 102 4,386 -
Massive 59 2,974 -
GoEmo 27 5,940 -
GCD $K$ $|D^{test}|$ $|D^{u}|$
Bank77 77 3,080 2,700
Clinc150 150 2,250 5,400
Stackoverflow 20 1,000 5,400
Table 1: Dataset Statistics in two scenarios. Emerging data scenario: we use the same settings as Zhang et al. (2023); Lin et al. (2025). GCD scenario: for fair comparison, we evaluate on the same test sets as De Raedt et al. (2023); Liang et al. (2024a); Zou et al. (2025b), but we only use $30\%$ of the original training set, which contains unlabeled data only. Further details are provided in the Appendix A.
5.1 Datasets and models
We use different topics of datasets for our experiments, including the intent datasets Bank77 (Casanueva et al., 2020), CLINC150 (Larson et al., 2019), Mtop (Li et al., 2020), Massive (Fitzgerald et al., 2023); the emotion dataset GoEmo (Demszky et al., 2020), and the community QA dataset Stackoverflow (Xu et al., 2015) for our experiments. Statistics are provided in Table 1.
For the emerging data scenario, we use the embedders e5-large (Wang et al., 2022) and instructor-large (Su et al., 2022) to compare with previous studies (Zhang et al., 2023; Feng et al., 2024). We also experiment with bert-base-uncased (Devlin et al., 2018), all-MiniLM-L6-v2 (Reimers and Gurevych, 2019). For the GCD scenario, we use all-MiniLM-L6-v2 (22.7M paramters) as the embedder, since most prior studies (An et al., 2024; Zou et al., 2025b) fine-tune bert-base (110M paramters). Using all-MiniLM-L6-v2 allows us to minimize the impact of model size on our experiments.
For the LLM experiments, the main results are obtained using gemma-2-9b-it (Rivière et al., 2024). Llama3.1-8B-Instruct AI@Meta (2024) and Qwen/Qwen3-8B (Team, 2025) are used for hyperparameter search or further analysis. An A100 GPU or an H100 GPU were used throughout the experiments.
5.2 Prompts used
Prompt for LLM
The prompts we used are shown in Table 2. It consists of three parts: task instruction, answer format, and task.
CLINC150, Bank77, Mtop, and Massive
Task Instructions:
Compare each Candidate Utterance to the Target Utterance. Select only Candidates with the same intent as the Target Utterance. Intent refers to the request or the purpose the user wants to achieve.
GoEmo
Task Instructions:
Compare each Candidate Utterance to the Target Utterance. Select only Candidates with the same emotion as the Target Utterance.
Stackoverflow
Task Instructions:
Compare each Candidate Question to the Target Question. Select only Candidates with the same main programming framework, language, tool, or concept as the Target Question.
Answer Format:
Only provide the final selection of Candidate Utterances by listing their numbers if they match the Target Utterance intent or request. 1. If Candidates 3, 4, and 9 match, write: The Candidate utterance number(s): 3, 4, 9 2. If no Candidates match, write: The Candidate utterance number(s): none Note: Stick to the answer format and avoid providing extra explanations.
Task:
Target Utterance: {sentence 1}
Candidate Utterances:
1. {sentence 1}
m. {sentence m}
Table 2: Task Instructions Prompt. Note: The boldface used here is for readability; it is not used in the prompt. If the dataset is Stackoverflow, we use ‘question’ instead of ‘utterance’ in the Answer format.
5.3 Hyperparameter setting
The main hyperparameter in our approach is the initial dimension (number of representatives) of the sparse vector $d$, which we set to 2,048 based on validation results on two external datasets: ArxivS2S and Reddit (Muennighoff et al., 2023). We use external datasets to prevent overfitting on a specific dataset. See details in Appendix B. Qwen3-8B (as LLM) and All-MiniLM-L6-v2 (as embedder) are used in the hyperparameter search.
5.4 Evaluation
The emerging and GCD scenarios are evaluated under two conditions: with HDBSCAN (McInnes et al., 2017) and K-means (Lloyd, 1982).
To reflect real-world settings with unknown $K$, we apply HDBSCAN (McInnes et al., 2017) to partition $D^{test}$ into $\hat{K}$ clusters. We set min_samples = 1 to reduce noise. The noise points are assigned to the nearest cluster centroid. We perform an incremental grid search over min_cluster_size values of $10,15,\ldots,50$, stopping early if the silhouette score does not improve (Rousseeuw, 1987; Gung et al., 2023).
K-means clustering is the dominant evaluation practice because many studies assume a known cluster number $K$ and incorporate it in their method pipelines Zhang et al. (2023); De Raedt et al. (2023); Liang et al. (2024b); Feng et al. (2024); Zou et al. (2025b). Following their convention, the number of clusters for K-means is set to match the number of ground-truth cluster number $K$.
To reduce cost, we follow (Zhang et al., 2023) and perform a single pipeline run per dataset using a fixed seed to obtain a sparse vector for each text. Most steps are deterministic. The main source of variability comes from splitting candidate examples across multiple LLM calls (e.g., three calls with 10 examples instead of one with 30), done only to save time and GPU memory. LLM outputs are generated via greedy decoding, without sampling. For evaluation, K-means is run over 5 different seeds and averaged. HDBSCAN, which is near-deterministic, is run once. After we derive the clusters, we apply standard clustering metrics for evaluation. These metrics include normalized mutual information (NMI) and clustering accuracy (Acc) (Rand, 1971; Meilă, 2007; Huang et al., 2014; Gung et al., 2023).
5.5 Baselines
Emerging Data Scenario
We compare against SCCL (Zhang et al., 2021), ClusterLLM, (Zhang et al., 2023), IDAS De Raedt et al. (2023), LLMedgeRefine (Feng et al., 2024), and SPILL (Lin et al., 2025). Table 3 shows that our method makes the fewest assumptions and has the lowest requirements. SPILL is the only method under the same assumptions. However, we still include ClusterLLM, IDAS, and LLMedgeRefine, which uses additional information – the known $K$ – in its development pipeline, to provide a more comprehensive comparison.
GCD Scenario
We compare against IDAS De Raedt et al. (2023), Loop (An et al., 2024), ALUP Liang et al. (2024b), IntentGPT Rodriguez et al. (2024) and Glean Zou et al. (2025b). Note that IntentGPT and IDAS can also be implemented only using unlabeled data. Table 3 shows that our method again makes the fewest assumptions and has the lowest requirements.
Emerging data FT known $K$ Labeled data
TWIST (Ours) -
SCCL (Zhang et al., 2021) $\checkmark$ -
ClusterLLM (Zhang et al., 2023) $\checkmark$ $\checkmark$ -
LLMEdgeRefine (Feng et al., 2024) $\checkmark$ -
IDAS (De Raedt et al., 2023) $\checkmark$ -
SPILL (Lin et al., 2025) -
GCD
TWIST (Ours)
Loop (An et al., 2024) $\checkmark$ $\checkmark$ $\checkmark$
ALUP (Liang et al., 2024b) $\checkmark$ $\checkmark$ $\checkmark$
IntentGPT (Rodriguez et al., 2024)
Glean (Zou et al., 2025b) $\checkmark$ $\checkmark$ $\checkmark$
Table 3: Baselines settings. FT: Finetuning. known $K$: Assuming known $K$ during the method pipeline development. Labeled data: Labeled data is required.
6 Results
6.1 Main results
Emerging data scenario
Table 4 shows the HDBSCAN results. Our approach substantially improves performance compared to directly using the embedder. [^1] The estimated number of clusters, $\hat{K}$, is provided in Appendix D. Moreover, the results from BERT (non-contrastive) also improve with our approach. For the K-means results, Table 5 shows that our approach generally outperforms IDAS and ClusterLLM on both metrics across different embedders, and achieves comparable results to LLMEdgeRefine. This indicates that, even without assuming prior knowledge of $K$ during method development – as the compared methods do – our approach remains effective and better reflects real-world practice. While SPILL achieves slightly higher NMI, our method attains substantially higher ACC.
Lastly, we found that the HDBSCAN results overall are similar to those of K-means, indicating that our method remains robust even without knowing $K$.
Bank 77 Clinc150 Mtop Massive GoEmo Average
NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc
Bert
Plain 0.84 1.69 73.03 50.87 63.95 40.97 2.86 8.18 1.35 12.05 28.41 22.75
TWIST (Ours) 76.33 60.23 93.23 85.91 73.86 47.47 72.02 62.32 21.01 14.19 67.29 54.02
MiniLM
Plain 78.78 60.39 86.79 67.71 67.30 55.91 62.55 48.15 11.31 18.62 61.35 50.16
TWIST (Ours) 85.61 75.29 93.86 86.44 74.52 50.36 76.67 67.51 21.63 23.27 70.46 60.57
E5
Plain 78.68 62.18 87.57 65.62 63.45 53.79 1.15 7.74 15.43 17.70 49.26 41.41
TWIST (Ours) 84.86 72.24 93.41 86.89 72.53 44.32 75.85 65.12 27.77 29.03 70.88 59.52
Instructor
Plain 82.50 67.34 92.29 79.27 67.63 54.31 67.99 53.54 0.25 11.48 62.13 53.19
TWIST (Ours) 85.63 73.77 93.64 84.62 74.06 46.81 76.52 64.07 31.64 33.65 72.30 60.58
Table 4: Five-benchmark results (Emerging Data Scenario, HDBSCAN). Gemma is used in TWIST. Plain: direct embedding clustering. Bold numbers: best within the embedder.
Bank 77 Clinc150 Mtop Massive GoEmo Average
NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc
Bert
Plain 50.72 (0.42) 29.07 (0.85) 72.90 (0.60) 49.48 (0.83) 62.24 (0.46) 26.78 (0.32) 49.12 (0.64) 30.49 (1.00) 10.77 (0.46) 12.90 (0.37) 49.15 29.75
TWIST (Ours) 76.23 (0.14) 60.49 (0.72) 93.25 (0.07) 86.43 (0.65) 71.79 (0.20) 36.41 (0.20) 72.00 (0.13) 60.56 (0.61) 18.66 (0.10) 18.76 (0.29) 66.38 52.53
MiniLM
Plain 79.08 (0.80) 61.27 (1.36) 89.46 (0.19) 73.73 (0.95) 67.41 (0.53) 31.40 (1.62) 70.37 (0.62) 54.04 (1.28) 10.84 (0.16) 13.89 (0.63) 63.43 46.87
TWIST (Ours) 85.57 (0.06) 74.07 (0.71) 93.65 (0.07) 86.35 (0.34) 71.92 (0.22) 37.83 (0.72) 76.18 (0.26) 66.15 (0.58) 20.69 (0.09) 24.44 (0.37) 69.60 57.77
E5
Plain 77.86 (0.35) 60.88 (0.89) 91.09 (0.17) 75.70 (0.36) 71.17 (0.28) 33.04 (0.74) 71.70 (0.84) 53.97 (2.00) 21.13 (0.44) 21.93 (0.60) 66.59 49.10
SCCL 77.34 (0.62) 63.60 (1.37) 91.89 (0.49) 77.96 (1.78) 70.42 ((0.34) 33.82 ((1.07) 71.57 (0.89) 54.48 (1.80) 20.05 (0.63) 22.03 (0.69) 66.25 50.38
ClusterLLM 84.16 (0.36) 70.13 (1.34) 92.92 (0.29) 80.48 (0.93) 74.46 (0.11) 37.22 (1.18) 74.39 (0.21) 56.08 (1.01) 22.23 (0.17) 22.22 (1.15) 69.63 53.23
SPILL 83.56 (0.47) 70.25 (1.56) 92.93 (0.08) 83.18 (0.78) 71.77 (0.35) 36.83 (1.10) 75.40 (0.51) 60.28 (1.81) 25.14 (0.36) 24.82 (0.86) 69.76 55.07
TWIST (Ours) 85.09 (0.19) 74.23 (0.88) 93.31 (0.04) 85.30 (0.58) 71.57 (0.18) 37.32 (1.67) 75.72 (0.22) 62.93 (1.28) 26.58 (0.23) 33.20 (0.54) 70.45 58.60
Instructor
Plain 82.42 (0.68) 65.49 (2.41) 93.27 (0.32) 81.68 (1.31) 71.63 (0.39) 33.14 (0.52) 74.02 (0.54) 54.62 (2.06) 29.18 (0.76) 32.51 (1.51) 70.10 53.49
SCCL 81.77 (1.36) 65.48 (1.36) 92.94 (0.44) 80.85 (0.74) 73.52 (0.38) 34.28 (0.58) 73.90 ((0.36) 54.10 ((1.05) 30.54 (0.68) 34.33 (0.86) 70.53 53.81
ClusterLLM 85.15 (0.41) 71.20 (1.59) 94.00 (0.21) 83.80 (0.41) 73.83 (0.79) 35.04 (0.97) 77.64 (0.21) 60.69 (0.96) 23.89 (0.68) 26.75 (1.76) 70.90 55.50
LLMEdgeRefine - (0.00) - (0.00) 94.86 (0.00) 86.77 (0.00) 72.92 (0.00) 46.00 (0.00) 76.66 (0.00) 63.42 (0.00) 29.74 (0.00) 34.76 (0.00) - -
IDAS - (0.00) - (0.00) 92.35 (0.00) 81.36 (0.00) 72.31 (0.00) 37.30 (0.00) 75.74 (0.00) 63.01 (0.00) 25.57 (0.00) 30.61 (0.00) - -
SPILL 85.01 (0.29) 71.05 (0.83) 93.77 (0.35) 85.14 (1.04) 72.65 (0.32) 37.11 (1.62) 77.62 (0.46) 62.42 (2.06) 32.78 (0.52) 33.21 (2.33) 72.37 57.79
TWIST (Ours) 85.46 (0.18) 74.66 (1.01) 93.92 (0.09) 85.62 (0.71) 72.34 (0.28) 38.40 (0.81) 76.43 (0.03) 63.17 (0.25) 30.62 (0.24) 37.04 (1.55) 71.75 59.78
Table 5: Five-benchmark results (Emerging Data Scenario, K-means). Averages over 5 runs (±SD). Gemma is used in TWIST. Plain: direct embedding clustering. Bold numbers: best within the embedder. SCCL/ClusterLLM/LLMEdgeRefine cited directly from prior work. SCCL results are cited from ClusterLLM; IDAS results are cited from LLMedgeRefine.
GCD scenario
Table 6 shows the HDBSCAN results for GCD. Our method consistently improves the performance of the existing embedder. Table 7 shows the K-means results for GCD. Our approach is on average the best in the unlabeled setting, without any known categories. It even outperforms baselines that know $10\%$ of the categories, and is comparable to other baselines that know $25\%$ of the categories (or up to $75\%$ in the case of IntentGPT). These strong results, together with the fewer assumptions required by our method, highlight its effectiveness and efficiency. Notably, BANK77 and CLINC150 perform similarly with both clustering algorithms, while Stackoveflow shows a noticeable drop.
Bank 77 Clinc150 StackO Average
KCR Method NMI Acc NMI Acc NMI Acc NMI Acc
- Plain 78.78 60.39 83.02 51.82 66.26 57.20 76.02 56.47
$\mathbf{0\%}$ TWIST 87.43 75.16 95.19 89.07 78.54 71.70 87.05 78.64
Table 6: Three-benchmark results (GCD scenario, HDBSCAN) Gemma is used in TWIST. Bold numbers: best. KCR ratio of known category. All-MiniLM-L6-v2 used as backbone embedder. Plain: direct embedding clustering. StackO: StackOverflow dataset.
Bank 77 Clinc150 StackO Average
KCR Method NMI Acc NMI Acc NMI Acc NMI Acc
- Plain 79.08 61.27 89.46 73.73 72.49 72.68 80.34 69.08
$\mathbf{0\%}$
IDAS 82.84 67.43 93.82 85.48 81.26 83.82 85.97 78.91
IntentGPT 81.42 64.22 94.35 83.20 - - - -
TWIST (Ours) 87.67 77.10 95.21 89.07 81.69 82.56 88.19 82.91
$\mathbf{10\%}$
Loop 79.14 64.97 93.52 84.89 75.98 80.50 82.88 76.79
GLEAN 82.23 67.99 95.21 88.71 79.67 82.40 85.70 79.70
$\mathbf{25\%}$
ALUP 84.06 74.61 94.84 88.40 76.58 82.20 85.16 81.74
Loop 83.37 71.40 94.38 86.58 79.10 82.20 85.62 80.06
GLEAN 85.62 76.98 96.27 91.51 80.90 84.10 87.60 84.20
$\mathbf{75\%}$ IntentGPT 85.94 77.21 96.06 88.76 - - - -
Table 7: Three-benchmark results (GCD scenario, K-means). Averages over 5 runs with Gemma. KCR: known categories ratio. All-MiniLM-L6-v2 used as backbone embedder in TWIST. Loop results are cited from GLEAN. Plain: direct embedding clustering.
6.2 Further analysis
Effect of Different LLMs
Our main results were obtained with gemma-2-9b-it. To investigate whether using a different LLM affects performance, Figure 2 shows the results obtained with all-MiniLM-L6-v2 as the embedder and three different LLMs (see Appendix E for details). Overall, the differences between Qwen3-8B and gemma-2-9b-it are similar but an obvious drop with Llama3.1-8B-Instruct. Because Qwen3-8B achieves comparable results, we use it along with all-MiniLM-L6-v2 for remaining analyses to reduce computational cost.
Figure 2: ACC and NMI across all test datasets (Emerging Data and GCD) for three LLMs. Results for K-means are averaged over 5 runs, HDBSCAN from a single run.
Effect of the Number of Closest Neighbors m and Iterative Updates
We use $m=30$ across all the experiment, but this raises the question of whether different values of $m$ would affect performance. Figure 3(a) shows that using too few neighbors leads to lower results. To investigate the effect of iterative updates, Figure 3(b) shows that in subsequent iterations, the results substantially improve, especially after the first iteration. Since most sparse vectors converge within 5 iterations, there are only minimal changes afterward: We find that most datasets have 90% convergence ratio by the 5th iteration (see Appendix F).
Figure 3: Comparison of (a) varying negihbor $m$ and (b) iterative updates on ACC and NMI across all test datasets (Emerging Data and GCD).
A scalable implementation of our method
Our representation relies on LLM guidance for each text, which can be computationally expensive when the dataset is extremely large. To assess scalability, we additionally explore clustering over the full dataset. The goal is to implement our approach on a subset of the texts and then use a model to generalize to the remaining texts (a form of knowledge distillation), without requiring further LLM computation. Specifically, we train a lightweight MLP ($\sim$10M parameters) to map pre-trained embeddings to sparse vectors. See Appendix G for experiment details. Note that we do not use any labeled data. Table 8 shows that our approach still improves the original embeddings by a large margin, which proves the scalability of our method, offering substantial savings in time and computational resources.
7 Conclusions
We propose TWIST (Short Text Clustering through Iterative Sparse Vector Updating with LLMs), an intuitive, domain-adaptive, label-free, and training-free method for short text clustering. Compared with other SOTA methods, TWIST is built under minimal assumptions and requirements. TWIST achieves better or similar results to state-of-the-art approaches without assuming prior knowledge of clusters or labels, making it more aligned with real-world scenarios. Experiments on diverse datasets and smaller LLMs show that our method is robust in low-resource settings, and also scalable to large datasets. In the future, we plan to expand our approach to clustering in specialized domains and explore ways to incorporate human feedback into our method pipeline.
Bank 77 Clinc150 StackOverf Average
Method NMI Acc NMI Acc NMI Acc NMI Acc
Kmeans
Plain 76.31 59.41 88.99 73.99 68.42 65.93 77.91 66.44
TWIST (Ours) 83.80 73.58 92.78 87.27 72.98 76.90 83.19 79.25
HDBSCAN
Plain 76.15 60.27 84.31 62.84 63.01 61.80 74.49 61.64
TWIST (Ours) 83.39 71.79 92.55 86.04 71.08 69.58 82.34 75.80
Table 8: Clustering results on the entire dataset using knowledge distillation. Averages over 5 runs for K-means. Plain: Directly embedding the test dataset with All-MiniLM-L6-v2.
Acknowledgments
This publication is part of the project LESSEN with project number NWA.1389.20.183 of the research program NWA ORC 2020/21 which is (partly) financed by the Dutch Research Council (NWO).
Limitations
Language. Like most of the prior work, we only focus on English utterance datasets. This is relevant because most of benchmark datasets are in English. Our method is completely data-driven and not dependent on labeled data and it is therefore applicable to other languages than English.
LLM capability. Our method relies on the abilities of LLMs. Although we empirically show that small LLMs can achieve good results, this is not guaranteed when the dataset comes from a rare domain.
Embedder capability. Although we empirically show that our approach generally improves performance compared to directly using the embedder for clustering, if the embeddings provide a poor initial result, the improvement might be limited.
Appendix A Comparison of Data Splits in GCD scenario: Our Method vs. Prior Work
Table 9 shows the train–validation–test splits used in Loop (An et al., 2024), ALUP (Liang et al., 2024b) and GLEAN Zou et al. (2025b) for the three benchmark datasets. In their setting, a specified ratio of all categories is randomly selected as known categories, referred to as the known category ratio (KCR). For each known category, 10% of the data is chosen to form the labeled dataset [^2] IntentGPT (Rodriguez et al., 2024) mention they use fewer than 10% for each known category, but they do not specify the exact percentage., while the remaining samples constitute the unlabeled dataset. Together, these two subsets form the training dataset. In our experiments, we do not use the validation set and use only 30% of the training dataset, restricted to the unlabeled setting with no access to labeled data. Evaluation is conducted on the same test split.
They train an embedder using the training split and optimizes it based on the labeled validation set, with the test set used solely for evaluation. In contrast, our method performs test-time embedding construction, so it requires to access test data, similar to IDAS De Raedt et al. (2023) and IntentGPT Rodriguez et al. (2024), using all available text data ($D=D^{test}\cup D^{u}$) to construct embeddings.
Since our clustering is unsupervised, no label information is used. In practice, clustering methods require access to the text to be clustered in order to construct representations, and our approach naturally uses all available texts.
# clusters # train # valid # test # total
Bank77 77 9,003 1,000 3,080 13,083
Clinc150 150 18,000 2,250 2,250 22,500
StackO 20 18,000 1,000 1,000 20,500
Table 9: Train–Validation–Test splits from previous studies (An et al., 2024; Liang et al., 2024b; Zou et al., 2025b). ‘StackO’ is the StackOverflow dataset
Appendix B Selection of $d$
Table 10 shows the two datasets that we used for selecting the dimensionality of the sparse vector $d$. Figure 4 shows that $1024$, $2048$, and $\min\{4096,N\}$ give similar performance for both K-means and HDBSCAN. We selected value $d=2048$ and used this value across all datasets and experiments.
[…]
ArxivS2S 93 3,674
Reddit 50 3,217
Table 10: External datasets used for selecting the dimensionality hyperparameter $d$
Figure 4: The average ACC and NMI across two datasets for different $d$. Results for K-means are averaged over 5 runs, while HDBSCAN is shown from a single run. $4096$ here means $\min\{4096,N\}$ because the dataset size can be smaller.
Appendix C Prompts for Instructor embedder
Our method is built on top of existing embedders. For instructor-large, a prompt can be added to generate task-specific embeddings. We follow previous research (Su et al., 2022; Liang and Liao, 2023; Lin et al., 2025) and use mostly the same prompts. The prompts we used for the different datasets are listed in Table 11.
Instructor:
Bank77: Represent the bank purpose for retrieval:
Mtop and Stackoverflow: Represent the sentence for retrieval:
Clinc150 and Massive: Represent the sentence for retrieving the purpose:
GoEmo: Represent an emotion sentence for retrieval:
Table 11: The prompts used in Instructor.
Appendix D Estimated Number of Clusters $\hat{K}$ by HDBSCAN
Table 12 and Table 13 show the estimated number of clusters, $\hat{K}$, for the two scenarios. The results indicate that the refined embeddings produced by our method are generally closer to the ground truth than the pre-trained (Plain) embeddings. An exception is the Bank77 dataset.
Bank 77 Clinc150 Mtop Massive GoEmo
Bert
Plain 2 111 62 2 2
TWIST 73 154 80 57 63
MiniLM
Plain 78 139 17 27 11
TWIST 68 150 62 56 44
E5
Plain 73 134 9 2 3
TWIST 70 142 82 57 58
Instructor
Plain 71 144 18 29 2
TWIST 84 148 99 58 57
True $K$ 77 150 102 59 27
Table 12: Emerging scenario: Estimated Number of Clusters $\hat{K}$ by HDBSCAN. gemma-2-9b-it used in TWIST.
Bank 77 Clinc150 Stackoverflow
Plain 78 82 13
TWIST 84 142 30
True $K$ 77 150 20
Table 13: GCD scenario: Estimated Number of Clusters $\hat{K}$ by HDBSCAN. All-MiniLM-L6-v2 used as backbone embedder. gemma-2-9b-it was used in TWIST.
Bank 77 Clinc150 Mtop Massive GoEmo Bank 77GCD Clinc150GCD StackOGCD Average
NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc
Kmeans
Gemma 85.57 74.07 93.65 86.35 71.92 37.83 76.18 66.15 20.69 24.44 87.67 77.10 95.21 89.07 81.69 82.56 76.57 67.20
Qwen 85.13 73.18 93.59 86.96 72.23 39.03 74.55 63.64 20.12 22.80 87.04 75.78 94.80 88.21 80.69 82.32 76.02 66.49
Llama 82.46 69.81 91.03 82.21 70.08 36.21 72.47 60.47 17.60 19.67 84.34 70.91 93.83 86.27 75.69 78.68 73.44 63.03
HCBSCAN
Gemma 85.61 75.29 93.86 86.44 74.52 50.36 76.67 67.51 21.63 23.27 87.43 75.16 95.19 89.07 78.54 71.70 76.68 67.35
Qwen 85.04 73.90 93.56 86.42 74.42 50.11 74.39 62.36 22.10 19.81 86.81 75.29 94.78 88.18 78.07 71.50 76.15 65.95
Llama 82.45 68.70 91.00 78.64 72.80 51.12 73.19 63.84 18.53 16.84 84.16 69.35 93.84 84.18 74.78 72.20 73.84 63.11
Table 14: Further analysis of results using different LLMs.
Appendix E Results by Different LLMs
We report the detailed results of all datasets by different LLMs in Table 14. Qwen and Gemma consistently outperform LLaMA. In addition, the average differences between K-means and HDBSCAN are minimal across both metrics for each LLM.
Appendix F Percentage of texts converged at iteration $t$
Figure 5 shows ratio of text convergence. It shows most datasets have 90% convergence ratio by the 5th iteration, and by the 7th iteration, all datasets achieve a 95% convergence ratio.
Figure 5: Ratio of texts converged at iteration $t$. All-MiniLM-L6-v2 (as embedder) and Qwen3-8B (as LLM) were used.
Appendix G Details of the full dataset clustering
Datasets Used with LLM Guidance
Table 15 shows the number of examples we use with the LLM. The number of examples with LLM guidance is the same as in Table 1, i.e. $|D^{test}|+|D^{u}|$, so we do not need to re-run all experiments. The examples account for 29.6% to 44.2% of datasets. Examples with LLM guidance are used to train the MLP, since their sparse vectors have been derived. The pre-trained embeddings from all-MiniLM-L6-v2 serve as input, and the LLM-generated sparse vectors serve as target output for supervised learning with an MLP. 20% of the examples are used as a validation set to tune the hyperparameters.
MLP Architecture and Hyperparameter Search
We build an MLP for each dataset. We only tune the hidden layer dimensionality using the validation set, with search over $\{512,1024,2048\}$.
Clustering of the entire dataset with MLP output
We first encode all examples into pre-trained embeddings using all-MiniLM-L6-v2, then use the trained MLP to derive sparse versions of these embeddings. These sparse vectors are used to cluster all examples.
Dataset $K$ Total LLM No-LLM LLM ratio
Bank77 77 13,083 5,780 7,303 44.2%
Clinc150 150 22,500 7,650 14,850 29.6%
Stackoverflow 20 20,000 6,400 13,600 32.0%
Table 15: Dataset Statistics.
Evaluation
For HDBSCAN, we set min_samples = 1 to reduce noise. The noise points are assigned to the nearest cluster centroid. We perform an incremental grid search over min_cluster_size values of $50,75,100,125,\ldots,250$ – five times the original range – stopping early if the silhouette score does not improve. For K-means, we set the number of clusters equal to the ground-truth cluster count $K$.

### v2

@@
-TWIST: Training-free and Label-free Short Text Clustering through Iterative Vector Updating with LLMs
+LLMs Enable Bag-of-Texts Representations for Short-Text Clustering
 Abstract
-In this paper, we propose a training-free and label-free method for short text clustering that can be used on top of any existing embedder.
-In the context of customer-facing chatbots, companies are dealing with large amounts of user utterances that need to be clustered according to their intent.
-In these commercial settings, no labeled data is typically available, and the number of clusters is not known.
-Our method is based on iterative vector updating: it constructs sparse vectors based on representative texts, and then iteratively refines them through LLM guidance.
-Our method achieves comparable or superior results to state-of-the-art methods that use contrastive learning, but without assuming prior knowledge of clusters or labels.
+In this paper, we propose a training-free method for unsupervised short text clustering that relies less on careful selection of embedders than other methods.
+In customer-facing chatbots, companies are dealing with large amounts of user utterances that need to be clustered according to their intent.
+In these settings, no labeled data is typically available, and the number of clusters is not known.
+Recent approaches to short-text clustering in label-free settings incorporate LLM output to refine existing embeddings.
+While LLMs can identify similar texts effectively, the resulting similarities may not be directly represented by distances in the dense vector space, as they depend on the original embedding.
+We therefore propose a method for transforming LLM judgments directly into a bag-of-texts representation in which texts are initialized to be equidistant, without assuming any prior distance relationships.
+Our method achieves comparable or superior results to state-of-the-art methods, but without embeddings optimization or assuming prior knowledge of clusters or labels.
 Experiments on diverse datasets and smaller LLMs show that our method is model agnostic and can be applied to any embedder, with relatively small LLMs, and different clustering methods.
-We also show that our method scales to large datasets, reducing the computational cost of the LLM.
-These low-resource, adaptable settings and the scalability of our method make it more aligned with real-world scenarios than existing clustering methods.
-Our source code is available here: https://anonymous.4open.science/r/sparse_vector-F4C9/README.md
+We also show how our method scales to large datasets, reducing the computational cost of the LLM use.
+The flexibility and scalability of our method make it more aligned with real-world training-free scenarios than existing clustering methods.
+Our source code is available here: https://anonymous.4open.science/r/BoT_vector-2E0C/README.md
 1 Introduction
+Figure 1: illustration of our proposed method.
+Left: traditional approach using the LLM output to refine the embeddings space.
+Right: our approach using the LLM output to directly construct a bag-of-texts space.
 Short text clustering is an NLP task that automatically groups unlabeled data into clusters and is widely used in practical applications for text classification and new category detection.
@@
 Other downstream tasks, such as mining topics and opinions from online platforms, also rely on effective clustering (Wu et al., 2024).
-The clustering task is often approached as a Generalized Category Discovery (GCD) scenario (Vaze et al., 2022), where limited labeled data and substantial unlabeled data are used to cluster all unlabeled samples, which include both known and unknown classes.
+The short-text clustering task is often approached as a Generalized Category Discovery (GCD) scenario (Vaze et al., 2022), where limited labeled data and substantial unlabeled data are used to cluster all unlabeled samples, which include both known and unknown classes.
 Recently, other studies have explored an emerging data scenario, where no labels are provided at all (De Raedt et al., 2023; Viswanathan et al., 2024b; Zhang et al., 2023; Lin et al., 2025).
 This scenario reduces reliance on human annotation and shows greater flexibility.
-A common approach for the emerging data scenarios is to train an embedder with contrastive learning to improve text representations (Zhang et al., 2021).
-With the advancement of LLMs, training-free approaches have also been proposed (De Raedt et al., 2023).
-The emerging data scenario has attracted increasing attention in recent research (Zhang et al., 2023; Feng et al., 2024; Lin et al., 2025).
-Despite the progress of prior work, these label-free approaches still face certain limitations and leave room for further improvement.
-First, methods involving LLM feedback often assume known number of clusters during method development (Zhang et al., 2023; Feng et al., 2024), which is difficult to estimate in real-world settings.
-Second, training-free methods have mainly been explored with large, proprietary models (Viswanathan et al., 2024b; De Raedt et al., 2023), which can be costly.
-Third, the scalability of these methods remains a concern.
-Motivated by these challenges, we propose a training-free approach that requires only a smaller LLM, alleviates the known-cluster assumption, and is scalable.
-The core idea of our method is to construct new representations (referred to as sparse vectors in this paper) such that texts belonging to the same cluster become more similar.
-Our method consists of two stages: we first construct sparse representations based on representative texts, and then iteratively refine them by combining LLM guidance with pre-trained embeddings.
-The contributions of our method are as follows:
-  • Inclusive and easy to optimize: Builds on existing embedders and requires only some hyperparameter tuning.
-  • Flexible: Does not assume a known number of clusters, nor require labeled data, contrastive learning on the task specific datasets, or fine-tuning, enabling easier adaptation to new domains.
-  • Low-resource: Even with small LLMs and embedders, the results are already competitive with methods that rely on labeled data or fine-tuning.
-  • Scalable: Does not require LLM feedback for every text, unlike many previous studies.
+Regardless the setting, the classic approach is to train an embedder with contrastive learning (Zhang et al., 2021).
+In recent years, most methods have been designed to use LLMs to guide the optimization (Zhang et al., 2023; Zou et al., 2025b).
+Many LLM-guided training-free approaches have been proposed as well to enable easier domain adaptation and reduce the need for fine-tuning (De Raedt et al., 2023; Viswanathan et al., 2024b; Lin et al., 2025).
+Although these training-free approaches lead to substantial improvement, they only partially capture LLM-derived similarities and remain heavily dependent on the original embeddings.
+The resulting representation may still not reflect LLM preference, despite this being a strong signal.
+Hence our objective: to construct text representations that directly incorporate LLM similarity estimations into the vector space.
+The core idea of our method (Figure 1) is to construct new representations (referred to as bag-of-texts vectors) such that texts belonging to the same cluster get more similar representations.
+Our method consists of two stages: we first construct bag-of-texts (BoT) representations based on representative texts, and then iteratively refine them by LLM guidance.
+In summary, the contributions of our method are as follows:
+  • A Strategy for Encoding LLM Preferences Unlike prior work that refines existing embeddings, we introduce a Bag-of-Texts (BoT) representation for encoding LLM outputs.
+  • Flexible: Our method can be added to any embedder and does not require fine-tuning, enabling easier adaptation to new domains.
+  • Low-resource: Our method reduces the need for human labeling, can be implemented using relatively small, zero-shot models, and achieves results competitive with labeled or fine-tuned approaches.
 2 Related Work
@@
 In practice, emerging data may show topic drift.
-Recently, with the advancement of LLMs, many studies have explored the emerging data scenario (Zhang et al., 2023; Viswanathan et al., 2024b; Lin et al., 2025), where no labels are available.
-We follow this unsupervised scenario.
-Boosting Unsupervised Short Text Clustering with LLM-Guided Contrastive Learning
-Contrastive learning is commonly used in unsupervised short text clustering (Zhang et al., 2021, 2023; Liang et al., 2024a).
+Recently, with the advancement of LLMs, many studies have explored the emerging data scenario (Zhang et al., 2023; Lin et al., 2025), where no labels are available.
+We follow this unsupervised scenario to support real world applications.
+LLM-Guided Contrastive Learning for Short Text Clustering
+With the advancement of LLM, recent work has begun to incorporate LLM feedback into the optimization process.
 Zhang et al.
-(2023) use LLMs to construct hard triplets to derive training pairs, which are then used to fine-tune a small embedder.
+(2023) use LLMs to construct hard triplets to derive training pairs.
 Liang et al.
-(2024a) use LLMs to refine texts by generating descriptors, then using small language models for contrastive learning to align the embeddings of texts and their descriptors.
+(2024b) identify samples in ambiguous regions and use LLMs to reassign them to new clusters, which are then used as training data.
+Zou et al.
+(2025b) proposed a unified framework that actively learns from diverse and quality-enhanced LLM feedback, including instance-level, category description, and uncertain-instance alignment.
 However, these methods either rely on knowing the number of clusters, proprietary models, or fine-tuning.
 Our method does not have these requirements.
-Training-Free Methods for Unsupervised Short Text Clustering
-Training-free approaches use a frozen pre-trained embedder: De Raedt et al.
+LLM-Guided Training-Free Approaches for Short Text Clustering
+Training-free approaches use a frozen pre-trained embedder and use LLMs in two ways.
+The first category is input text enrichment: De Raedt et al.
 (2023) select prototypical utterances and then use an LLM to generate a set of descriptive utterance labels for these prototypes.
-The labels and texts are then encoded to form better representations.
 Viswanathan et al.
-(2024a) use LLMs to improve texts by generating key phrases for each text, which are then separately encoded into embeddings and concatenated for clustering.
-Lin et al.
+(2024a) use LLMs for few-shot in-context learning to improve texts by generating key phrases for each text.
+The second category is embedding post-processing: Lin et al.
 (2025) use LLMs for selection rather than text enrichment.
 They first encode all of the texts, and then use the LLM to select similar pairs for each text.
-Each text embedding is then obtained by average pooling the embeddings of the text and its similar pairs.
-These methods have in common that they refine original embeddings to improve clustering.
+Each text embedding is then obtained by average pooling its similar pairs.
+These methods have in common that they refine original embeddings.
 In contrast, we construct a new vector using LLM feedback.
-Feng et al.
-(2024) use LLMs to refine edge points.
-However, the assumption of a known number of clusters can still be a core problem in real-world applications.
-3 Theoretic formulation and assumptions
+3 Bag-of-Texts Theoretical Justification
 3.1 Problem Definition
-In our work, we consider two variations of the short text clustering task: (i) the emerging data scenario and (ii) the GCD scenario.
-In the emerging data scenario (Lin et al., 2025), the only available resource is a collection of unlabeled texts $D^{test}\triangleq\{x_{i}\}_{i=1}^{N}$, where each $x_{i}\in D^{test}$ corresponds to a short text and $N$ is the total number of samples.
+The task is to partition $D^{test}$ into $K$ clusters.
+$K$ is treated as prior knowledge in most existing work (De Raedt et al., 2023; Zhang et al., 2023; Feng et al., 2024; Zou et al., 2025b), and many approaches incorporate $K$ when refining embeddings.
+Our proposed method does not make any assumption on the knowledge of $K$ thus can be implemented without knowing the number of clusters.
+For evaluation, however, we incorporate $K$ for fair comparison to previous methods.
+In line with prior work, we address two scenarios: (i) the emerging data scenario and (ii) the Generalized Category Discovery (GCD) scenario.
+In the emerging data scenario (Zhang et al., 2023; Lin et al., 2025), the only available resource is a collection of unlabeled texts $D^{test}\triangleq\{x_{i}\}_{i=1}^{N}$, where each $x_{i}\in D^{test}$ corresponds to a short text and $N$ is the size of the test dataset.
 The dataset is used as a whole for clustering and evaluation, without any train-test split.
-In contrast, the GCD scenario has access to additional data, typically including some labeled samples (Zou et al., 2025b).
-In this work, we adapt this setting by considering only an additional unlabeled dataset $D^{u}\triangleq\{x_{i}\}_{i=1}^{N_{u}}$, in line with the goal of low-resource.
-The GCD scenario is included for the purpose of testing whether extra data improves clustering and generalizability of our method.
-For both scenarios, the task is to partition $D^{test}$ into clusters $\{\hat{S}_{l}\}_{l=1}^{\hat{K}}$, where $\hat{K}$ is the number of predicted clusters.
-Note that $\sum_{l=1}^{\hat{K}}|\hat{S}_{l}|=N$.
-The goal is to ensure that the clustering results $\{\hat{S}_{l}\}_{l=1}^{\hat{K}}$ are as close as possible to the true partition $\{S_{l}\}_{l=1}^{K}$, where $K$ is the number of unique clusters in the ground truth.
-Importantly, unlike many existing studies (De Raedt et al., 2023; Feng et al., 2024; Zou et al., 2025b), which assume that $K$ is known in advance, we make no such assumption.
+In contrast, the GCD scenario typically has access to additional data, typically including some labeled examples (Liang et al., 2024b; Zou et al., 2025b).
+As opposed to prior work, to align the GCD scenario with the objective of low-resource settings, we restrict the availability of additional data to only an unlabeled dataset $D^{u}\triangleq\{x_{i}\}_{i=1}^{N_{u}}$.
 For notation convenience, we define the full dataset as $D\triangleq D^{test}\cup D^{u}$ to simplify the explanation of our proposed method. Note that $D=D^{test}$ in the emerging data scenario.
-3.2 Constructing the Sparse Vector
-Basic Assumption
-Considering the fact that texts belonging to the same cluster are similar to each other, we represent each text as a sparse bag-of-texts vector that directly encodes its similarity to other texts.
-Formally, for any $x_{i},x_{j}\in D$ with $i\neq j$:
-  • $x_{i}$ and $x_{j}$ are similar if $x_{i},x_{j}\in S_{k}$ for some cluster $S_{k}$.
-  • $x_{i}$ and $x_{j}$ are not similar if $x_{i}\in S_{k}$ and $x_{j}\in S_{l}$ with $k\neq l$.
-Construction of Bag-of-Texts
-Since similarity is determined by clusters, it is not necessary to use all texts in $D$ to represent sparse vectors; instead, a representative subset of texts can constitute our bag of texts, denoted as $D^{r}\subseteq D$, where $|D^{r}|\triangleq d$.
-The set of representative texts $D^{r}$ should contain at least one element from each cluster.
-That is, for any cluster $S\in\{S_{l}\}^{K}_{l=1}$, there exists at least one text $x\in D^{r}$ such that $x\in S$.
-Therefore, to cover all possible clusters, the minimum size of $d$ is K1K-1.
-Since $K$ is unknown, the most practical approach is to choose a relatively large $d$.
-For each $x\in D$, we define a $d$-diemnsional vector $\mathbf{z}$ assigned to the representative texts, where the $j$-th coordinate takes the value 1 if $x$ is similar to the $j$-th representative text, and 0 otherwise.
+3.2 Embedding Refinement via Bag-of-Texts
+Existing LLM-guided short-text clustering approaches often refine embeddings either directly using contrastive learning Zhang et al.
+(2023), indirectly by text enrichment De Raedt et al.
+(2023); Viswanathan et al.
+(2024b), or through embedding post-processing Lin et al.
+(2025).
+While contrastive optimization enforces inequality through gradient updates, it requires training on new samples, which makes it impractical for low-resource settings.
+Text enrichment approaches, on the other hand, can perform for low-resource scenarios but lack explicit guarantees for pulling similar texts closer and pushing dissimilar ones apart.
+Embedding post-processing approaches heavily depend on the embedder used.
+Our novel bag-of-texts representation combines the strengths of these approaches and directly enforces inequality through text-based LLM outputs, without requiring training data or expensive gradient-based updates of embedders.
+Our approach mimics contrastive learning by reconstructing the vector space using a bag-of-texts representation and updating this representation using LLM guidance.
+Using bag-of-texts representation, all texts start with equivalent distances in the vector space, and LLM output about similarity of texts to each other is directly translated to the vector representation, moving each text toward the correct region of the space; see Figure 1 for illustration.
+Formally, for a text $x$, let $l_{x,x^{+}}$ and $l_{x,x^{-}}$ denote similarities between $x$ and a positive and negative example, and $l^{{}^{\prime}}_{x,x^{+}}$ and $l^{{}^{\prime}}_{x,x^{-}}$ be the corresponding similarities after embedding refinement.
+The positive and negative examples are identified by a function (e.g., an LLM or a human), which determines whether the text $x$ belongs to the same cluster as example $x^{+}$ (positive) or to a different cluster from the example $x^{-}$ (negative).
+Ideally, our refinement approach should satisfy $l^{{}^{\prime}}_{x,x^{+}}>l^{{}^{\prime}}_{x,x^{-}}$.
+Below, we formally explain how our bag-of-texts representation achieves this goal.
+Assumption 1.
+We assume that a function (e.g., an LLM or a human) makes a selection based on the following rules.
+  • Similarity: For any texts $x_{i}$ and $x_{j}$ with $i\neq j$, $x_{i}$ and $x_{j}$ are similar if they belong to the same cluster $S_{k}$, and not similar if they belong to different clusters $S_{k}$ and $S_{h}$ with $k\neq h$.
+  • Distinct Clusters: Each text is assigned to exactly one cluster $S\in\mathcal{S}$.
+Note that $\mathcal{S}$ in Assumption 1 refers to the clusters defined by the function; there is no knowledge of a ground truth.
+We assume distinct clusters as is conventional in prior work (An et al., 2024; Lin et al., 2025).
+Definition 1 (Bag-of-Texts).
+Let $g(x,D)$ be a function that takes a text $x$ and a set of texts $D$, and outputs a bag-of-text representation $\mathbf{z}\in\mathbb{R}^{|D|}$.
+In this representation, the entry corresponding to each text in $D$ that belongs to the same cluster as $x$ is assigned a value of 1, while all other entries are 0.
 Proposition 1.
-Given the assumption, for any $x_{i},x_{j}\in D$ with $i\neq j$ that belong to the same cluster $S_{k}$, the corresponding sparse vectors $\mathbf{z}_{i}$ and $\mathbf{z}_{j}$ will be identical.
-The proposition is straightforward, so we omit the proof.
-Adaptation to real-world practice
-Although the proposition holds in the idealized setting, in practice similarity is typically determined using LLMs, which may introduce errors, and scanning the entire dataset for each $x$ is computationally expensive.
-To address these, we construct a sparse vector $\mathbf{z}$ in a way that is adaptable to real-world practice: for each $x\in D$, instead of assigning fixed value in $\{0,1\}$ based on representative texts, we propose an updating procedure that accounts for uncertainty.
-In addition, to address limited computational resources, we use an existing embedder to identify a relatively small set of $m$ similar candidates that are likely to belong to the same category, as empirically shown in previous research (Lin et al., 2025; Zou et al., 2025a).
-The details of the method will be discussed in the next section.
+Given Assumption 1, for any $x_{i},x_{j}\in D$ and their corresponding bag-of-texts representation $\mathbf{z}_{i},\mathbf{z}_{j}$, the function $g(x,D)$ guarantees that $l^{{}^{\prime}}_{x,x^{+}}>l^{{}^{\prime}}_{x,x^{-}}$.
+The proof of the proposition builds on Definition 1 that if two texts $x_{i},x_{j}\in D$ belong to the same cluster, their bag-of-texts representations $\mathbf{z}_{i}$ and $\mathbf{z}_{j}$ are identical.
+Considering cosine similarity as an example similarity function, $\text{cosine}(\mathbf{z}_{i},\mathbf{z}_{j})=1$ if they are identical, and $\text{cosine}(\mathbf{z}_{i},\mathbf{z}_{j})=0$ if they are dissimilar.
+3.3 Adaptation to Real-world Practice
+Although the proposition holds in the idealized setting, in practice the similarity between texts is typically determined using LLMs, which may introduce errors, and scanning the entire dataset for each $x$ is computationally expensive.
+We outline the modifications needed for practical implementation.
+The detailed analysis is in appendix A.
+Memory constraints.
+Large datasets lead to high-dimensional bag-of-text vectors.
+To address these issues, we construct bag-of-texts vectors using a representative subset $D^{r}\subseteq D$ with $|D^{r}|\triangleq d$ , rather than all texts.
+This dimensionality reduction is possible because it relies on the assumption that texts within the same cluster are similar; therefore, it is not necessary to use all texts in $D$ to represent bag-of-texts vectors.
+Instead, the bag-of-texts vectors of the remaining texts can be represented via the vectors of the selected representatives.
+Computation constraints.
+Selecting similar texts requires scanning the entire dataset for each text, which is computationally expensive with an LLM.
+To reduce computation, we use a pre-trained embedder to retrieve a small set of $m$ candidate texts that are likely to be similar, as supported by prior work (Lin et al., 2025; Zou et al., 2025a).
+Selection uncertainty.
+During automatic selection (e.g., by LLM) inconsistent preferences can occur and violate Assumption 1.
+For example, $(x_{i},x_{j})$ may be identified as belonging to the same cluster, $(x_{i},x_{l})$ as different, yet $(x_{j},x_{l})$ as the same.
+Also, in practice, each selection is made with some inherent confidence.
+To reduce these inconsistencies, we perform iterative update and assign the mean of the selected bag-of-texts vectors rather than a fixed value of 1.
 4 Computational Method
-Our method aims to construct sparse vectors that capture the similarity of texts within the same cluster.
+Our method aims to construct Bag-of-Texts (BoT) vectors that capture the similarity of texts within the same cluster.
 It consists of two stages.
-In the initial stage, we select representative texts, which are then used to construct a sparse vector for each text.
-In the second stage, we iteratively update each text’s sparse vector by considering its similar texts until convergence, illustrated in Figure 1.
-Unlike many prior studies that assume the number of clusters $K$ is known, we consider a more practical setting in which $K$ is unknown, and thus approximate the true number of clusters by the number of representative texts, $d$.
-This is less challenging than estimating $K$, as it only requires ensuring that $d$ is large enough to cover $K$ clusters.
-Figure 1: Construction of Sparse vectors.
-Initial stage: We first encode each text in $D$ using a pretrained embedder and partition them into $d$ clusters.
-We then select the medoid of each cluster as the representative text and construct one-hot vectors for these representatives.
-The remaining text vectors are obtained based on LLM selection.
-Iterative stage: We update all of the vectors iteratively until convergence.
-4.1 Initial stage: construction of sparse representation
-To build sparse vector representations, we first need to create a set of representative samples $D^{r}$.
-To this end, we first encode each text using an existing embedder $F$ to obtain pre-trained embeddings:
-$\mathbf{x}_{i}=F(x_{i}).$
-Next, we apply a clustering algorithm to partition these embeddings into $d$ clusters.
-We use agglomerative clustering (Murtagh and Legendre, 2014) as it produces deterministic results, which facilitates reproducibility.
-For each initial cluster , we derive the medoid.
-We refer to the set of medoids as the representative text, denoted as:
-$D^{r}\triangleq\{x_{c_{1}},\dots,x_{c_{d}}\}.$
-and let $\mathbf{X}^{r}$ be $D^{r}$ corresponding pre-trained embeddings, where $\mathbf{X}^{r}=\{\mathbf{x}_{i}:x_{i}\in D^{r}\}$.
-We then transform the representative texts into one-hot vectors, which serve as the basis vectors for the sparse representation.
-For the remaining texts $x_{j}\in D\setminus D^{r}$, we construct their sparse vectors using this basis.
-For each $x_{j}$, we compute the pairwise cosine distances between its embedding $\mathbf{x}_{j}$ and the medoid embeddings in $\mathbf{X}^{r}$.
-By sorting these distances, we select the top $m$ closest representative texts to $x_{j}$, denoted as $x_{j1},\dots,x_{jm}$.
-Then, we use an LLM to identify the similar text(s) to $x_{j}$ among these representative texts, denoted as $M_{j}\triangleq LLM(x,x_{j1},\dots,x_{jm})\in\mathcal{P}(\{x_{j1},\dots,x_{jm}\})\cup\{\text{None}\}$, where $\mathcal{P}(\cdot)$ represents the power set, i.e., all possible subsets of the candidate texts, and ‘None’ indicates that no representative text is selected.
-If the LLM output is not None, the sparse vector of $x_{j}$ is then derived by averaging the sparse vectors of the texts in $M_{j}$, denoted as $\mathbf{z}_{j}$.
-Instead of assigning a fixed value of 1 to the selected dimension, we average the vectors to capture how strongly each representative text contributes to the feature.
-We then apply L2 normalization to ensure a consistent scale for the iterative stage.
-If the LLM output is None, we make this text into a representative text itself and increase the dimensionality of all sparse vectors by one and construct its corresponding one-hot vectors.
-Note: if we set $d=N$, all texts will be representative texts and one-hot encoded.
-4.2 Iterative stage: sparse vectors refinements
-In this stage, we update the sparse vectors of all texts using the dynamic subset of $m$ candidates.
-At each iteration step $t$ (with the initial stage being $t=0$), for each $x_{j}\in D$, we first concatenate its pre-trained embedding and sparse vector: $\mathbf{u}_{j}^{t}\triangleq\mathbf{x}_{j}\oplus\mathbf{z}_{j}^{t}$.
-We then compute the pairwise cosine distances between $\mathbf{u}_{j}^{t}$ and the concatenated representations of the all texts $\mathbf{U}^{t}$.
-By sorting these distances, we select the top $m$ closest texts to $x_{j}$, denoted as $x^{t}_{j1},\dots,x^{t}_{jm}$.
-The concatenation ensures that the pre-trained embedding guides the sparse vectors.
-Also, the reason we do iterative updating instead of just once is to gain more stable sparse vectors.
-Next, we use an LLM to identify the subset of the texts most similar to $x_{j}$, denoted as $M^{t}_{j}\triangleq LLM(x_{j},x^{t}_{j1},\dots,x^{t}_{jm})\in\mathcal{P}(\{x^{t}_{j1},\dots,x^{t}_{jm}\})\cup\{\text{None}\}$.
-If the LLM output is None, we fix sparse vectors, i.e., $\mathbf{z}_{j}^{t+1}=\mathbf{z}_{j}^{t}$.
-Otherwise, we update the sparse vectors by averaging the sparse vector of the texts in $M^{t}_{j}$.
-We then apply L2 normalization, as explained previously.
-For $x_{j}$, the update process is ended if the sparse vector at iteration $t+1$ is almost identical to those at iteration $t$, as we define by a cosine similarity greater than 0.99.
-For efficiency, we set $m=30$ and cap the maximum number of iterations at 10.
+First, we select representative texts, which are then used to construct a BoT vector for each text.
+Second, we iteratively update each text’s BoT vector by considering its similar texts until convergence.
+4.1 Initial stage: construction of BoT representation
+Algorithm 1 Select Representative Texts
+0:$D$, $\mathbf{X}$, representative texts size $d$
+[…]
+[…]
+[…]
+3:for $j=1,\dots,d$ do
+4:  $\mathbf{x}^{r}_{j}=\arg\min_{\mathbf{x}\in C_{j}}\sum_{\mathbf{y}\in C_{j}}\|\mathbf{x}-\mathbf{y}\|_{2}$
+5:  $D^{r}\leftarrow D^{r}\cup\{x_{i}\in D:\mathbf{x}_{i}=\mathbf{x}^{r}_{j}\}$
+6:end for
+7: Initialize $\mathbf{Z}^{r}\leftarrow I_{d}$ {one-hot vectors for $D^{r}$}
+8:return Initial Representative Texts $D^{r}$, $\mathbf{Z}^{r}$
+Algorithm 2 Construct Bag-of-Texts Vectors
+[…]
+0: Bag-of-texts vectors $\mathbf{Z}\in\mathbb{R}^{N\times(d+d^{*})}$
+1: Initialize $\mathbf{z}_{j}\leftarrow\mathbf{0}$ for all $x_{j}\in D\setminus D^{r}$
+2:for each $x_{j}\in D\setminus D^{r}$ do
+3:  $M_{j}\leftarrow\{x_{j1},\dots,x_{jm}\}\in D^{r}$ from $\mathbf{X}$ as the top-$m$ by cosine similarity to $x_{j}$
+4:  $O_{j}\leftarrow\text{LLM}(M_{j})\in\mathcal{P}(M_{j})\cup\{\emptyset\}$ {# $\mathcal{P}(.)$: power set}
+5:  if $O_{j}\neq\emptyset$ then
+6:   $\mathbf{z}_{j}\leftarrow\text{MeanPooling}\{\mathbf{z}_{k}:x_{k}\in O_{j}\}$
+7:   $\mathbf{z}_{j}\leftarrow\frac{\mathbf{z}_{j}}{\|\mathbf{z}_{j}\|_{2}}$
+8:  end if
+9:end for
+[…]
+[…]
+12: Let $\mathbf{P}\leftarrow\{\mathbf{z}_{j}:\mathbf{z}_{j}=\mathbf{0}\}$; one-hot encode each $\mathbf{z}_{j}\in\mathbf{P}$ in new dimensions, and pad the first $d$ columns with zeros.
+13: Let $\mathbf{Q}\leftarrow\{\mathbf{z}_{j}:\mathbf{z}_{j}\neq\mathbf{0}\}$; pad $\mathbf{Q}$ and $\mathbf{Z}^{r}$ with $d^{*}$ zero columns
+14:return $\mathbf{Z}\leftarrow\text{Stack}(\mathbf{Z}^{r},\mathbf{P},\mathbf{Q})$, rows ordered according to $D$
+We first encode all texts $D$ with a pre-trained embedder to obtain embeddings $\mathbf{X}$.
+To select initial representative texts $D^{r}$, we perform agglomerative clustering (Murtagh and Legendre, 2014) to partition them into $d$ clusters, $d$ being a trivially large number (see Section 5.3).
+The medoid of each cluster is chosen as a representative and converted into a one-hot vector with initially equal spacing.
+See Algorithm 1 for specification.
+The remaining texts are positioned based on LLM preference.
+Specifically, for each text $x_{j}\in D\setminus D^{r}$, we retrieve the top $m$ representative texts from the pretrained embeddings $\mathbf{X}$ that are closest based on cosine similarity.
+An LLM is then used to select which representative(s) are similar.
+The bag-of-texts vector for each text is computed by averaging the vectors of its selected representatives.
+After processing all texts, those with no selected representatives are promoted to new representatives and added to $D^{r}$ (so that $|D^{r}|=d+d^{*}$, where $d^{*}$ is the number of such new representatives), and the dimensionality of the BoT vectors is expanded to match.
+See Algorithm 2 for specification.
+This procedure ensures that the initial BoT vectors are evenly spaced for the representative texts, while the vectors for the remaining texts encode LLM-informed relationships, avoiding the need to use the full dimensionality of all texts.
+4.2 Iterative stage: bag-of-texts vectors refinements
+After constructing the initial positions for all texts, we iteratively update the BoT vectors to better align their relative distances with LLM preferences.
+The update procedure follows the same approach as in Algorithm 2, with three key modifications.
+First, the update uses all texts as the pool for subset selection, since each text has a BoT vector.
+Second, for each text, we select a subset of $m$ candidates based on the concatenation of its pre-trained embedding and BoT vector, $\mathbf{U}^{t}\triangleq\mathbf{X}_{j}\oplus\mathbf{Z}_{j}^{t}$.
+This ensures that the pre-trained embedding initially guides the BoT vectors, which carry minimal information at the start.
+Third, the update is performed iteratively rather than just once, allowing the candidate set to vary dynamically as the BoT vectors are refined.
+The update for a text stops when its BoT vector changes negligibly, as measured by a cosine similarity greater than 0.99.
+We set $m=30$ and cap the maximum number of iterations at $T=10$ as budget.
+See Algorithm 3 in Appendix B for specification.
+Note that the pretrained embeddings $\mathbf{X}$ are used only for updating the Bot vectors; only the final BoT vectors $\mathbf{Z}^{T}$ will be used for clustering.
 5 Experimental Settings
-Emerging data $K$ $|D^{test}|$ $|D^{u}|$
+Emerging data scenario $K$ $|D^{test}|$ $|D^{u}|$
 Bank77 77 3,080 -
@@
 GoEmo 27 5,940 -
-GCD $K$ $|D^{test}|$ $|D^{u}|$
+GCD scenario $K$ $|D^{test}|$ $|D^{u}|$
 Bank77 77 3,080 2,700
@@
 (2024a); Zou et al.
-(2025b), but we only use $30\%$ of the original training set, which contains unlabeled data only.
-Further details are provided in the Appendix A.
-5.1 Datasets and models
-We use different topics of datasets for our experiments, including the intent datasets Bank77 (Casanueva et al., 2020), CLINC150 (Larson et al., 2019), Mtop (Li et al., 2020), Massive (Fitzgerald et al., 2023); the emotion dataset GoEmo (Demszky et al., 2020), and the community QA dataset Stackoverflow (Xu et al., 2015) for our experiments.
+(2025b), but we only use $30\%$ of the original unlabeled training set.
+Further details are provided in Appendix C.
+5.1 Datasets
+We use a variety of datasets to benchmark our method: the intent datasets Bank77 (Casanueva et al., 2020), CLINC150 (Larson et al., 2019), Mtop (Li et al., 2020), and Massive (Fitzgerald et al., 2023); the emotion dataset GoEmo (Demszky et al., 2020), and the community QA dataset Stackoverflow (Xu et al., 2015).
 Statistics are provided in Table 1.
-For the emerging data scenario, we use the embedders e5-large (Wang et al., 2022) and instructor-large (Su et al., 2022) to compare with previous studies (Zhang et al., 2023; Feng et al., 2024).
-We also experiment with bert-base-uncased (Devlin et al., 2018), all-MiniLM-L6-v2 (Reimers and Gurevych, 2019).
-For the GCD scenario, we use all-MiniLM-L6-v2 (22.7M paramters) as the embedder, since most prior studies (An et al., 2024; Zou et al., 2025b) fine-tune bert-base (110M paramters).
-Using all-MiniLM-L6-v2 allows us to minimize the impact of model size on our experiments.
+Among these, MTOP, Massive, and GoEmo have highly imbalanced clusters.
+5.2 Models
+To see how sensitive our approach is to the choice of embedder, we design experiments that incorporate multiple models.
+We experiment in the emerging data scenario using TF–IDF, bert-base-uncased (Devlin et al., 2018), all-MiniLM-L6-v2 (Reimers and Gurevych, 2019), and e5-large (Wang et al., 2022).
+To see how far our method is from semi-supervised approaches and the effect with additional data, in the GCD scenario, we use all-MiniLM-L6-v2 (22.7M parameters) as the embedder following prior work (Rodriguez et al., 2024).
 For the LLM experiments, the main results are obtained using gemma-2-9b-it (Rivière et al., 2024).
-Llama3.1-8B-Instruct AI@Meta (2024) and Qwen/Qwen3-8B (Team, 2025) are used for hyperparameter search or further analysis.
 An A100 GPU or an H100 GPU were used throughout the experiments.
-5.2 Prompts used
-Prompt for LLM
-The prompts we used are shown in Table 2. It consists of three parts: task instruction, answer format, and task.
+Prompts used
+The prompts we used for the LLM are shown in Appendix E.
+Following prior work (Zhang et al., 2023; De Raedt et al., 2023; Viswanathan et al., 2024a; Lin et al., 2025; Zou et al., 2025b), we keep the prompts simple, as is standard practice, to ensure a fair comparison and generalizability between models.
+5.3 Hyperparameter setting
+The main hyperparameter in our approach is the initial dimension $d$ (number of initial representatives).
+Although we can simply set the number of $d$ equal to the total number of examples $N$, this becomes memory-intensive when $N$ is extremely large.
+Following the approach in (Feng et al., 2024; Lin et al., 2025), we determine hyperparameters using external datasets to avoid biasing evaluation on the test sets.
+The search space starts from 512, setting a sufficiently large minimum to ensure that $d$ to include at least one representative from each cluster.
+We fix the BoT vector dimension to 1,024, based on validation results from two external datasets, including ArxivS2S and Reddit (Muennighoff et al., 2023).
+Empirically, we find that performance remains very similar between different values of $d$ (see Appendix D).
+5.4 Evaluation
+The Bag-of-Text vectors are evaluated under K-means.
+K-means clustering is a widely adopted evaluation practice in short text clustering Zhang et al.
+(2023); De Raedt et al.
+(2023); Viswanathan et al.
+(2024b); Liang et al.
+(2024b); Lin et al.
+(2025); Zou et al.
+(2025b).
+Following convention, the number of clusters for K-means is set to match the number of ground-truth cluster number $K$ during evaluation.
+K-means is run over 5 different seeds and averaged.
+After we derive the clusters, we apply standard clustering metrics for evaluation.
+These metrics include normalized mutual information (NMI) and clustering accuracy (Acc) (Rand, 1971; Meilă, 2007; Huang et al., 2014; Gung et al., 2023).
+5.5 Baselines
+We compare our method with SOTA representation learning baselines for text clustering.
+Table 2 summarizes these methods.
+For the emerging data scenario, we compare against IDAS De Raedt et al.
+(2023), ClusterLLM (Zhang et al., 2023), and SPILL (Lin et al., 2025).
+For the GCD Scenario we compare against IntentGPT Rodriguez et al.
+(2024).
+We also include comparisons with methods that require some labeled data, including LOOP An et al.
+(2024), ALUP Liang et al.
+(2024b) and Glean Zou et al.
+(2025b) to measure how far our method is from the performance of these semi-supervised approaches.
+Unsup LLM FT $\mathbf{K}$
+LeBoT (Ours) $\checkmark$ Gemma
+IDAS De Raedt et al. (2023) $\checkmark$ Gemma $\checkmark$ $\checkmark^{*}$
+ClusterLLM Zhang et al. (2023) $\checkmark$ GPT-* $\checkmark$ $\checkmark$
+IntentGPT Rodriguez et al. (2024) $\checkmark^{*}$ GPT-*
+ALUP (Liang et al., 2024b) GPT-* $\checkmark$ $\checkmark$
+LOOP An et al. (2024) GPT-* $\checkmark$ $\checkmark$
+SPILL Lin et al. (2025) $\checkmark$ Gemma
+Glean Zou et al. (2025b) GPT-* $\checkmark$ $\checkmark$
+Table 2: LLM-based representation learning for short-text clustering.
+Unsup: No access to labels.
+IntentGPT can be semi-supervised as well.
+LLM: Model used in the main results (Gemma = Gemma2-9b-it; GPT-* = GPT variants with unknown parameters).
+FT: Finetuning.
+$\mathbf{K}$: Incorporates the number of clusters $K$ into embedding refinement.
+IDAS can be applied either without knowing $K$ or knowing $K$
+6 Results
+6.1 Main results
+Emerging data scenario
+Table 3 reports the K-means results.
+Our approach consistently outperforms pretrained embeddings and on average performs better than all other baselines, including those that require fine-tuning (ClusterLLM).
+Furthermore, our approach maintains consistent performance across backbones, unlike SPILL and IDAS, which degrade more when the embedder changes.
+We also observe that with the stronger E5 embedder, which has the highest zero-shot performance, the gap between methods narrows.
+This is likely because E5 provides richer semantic representations and better separation in the embedding space, making it easier for both text-enrichment approaches (IDAS) and embedding post-processing methods (SPILL) to perform well.
+See Appendix F for detailed embedder analysis.
+Bank 77 Clinc150 Mtop Massive GoEmo Average
+NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc
+TF-IDF
+Plain 59.76 (0.31) 36.73 (0.99) 58.05 (0.43) 33.44 (1.11) 50.77 (0.84) 30.98 (1.19) 45.24 (1.55) 28.35 (1.61) 18.49 (1.65) 19.64 (1.25) 46.46 29.83
+IDAS 72.62 (0.28) 51.96 (0.78) 76.37 (0.68) 54.44 (1.04) 63.95 (0.11) 39.47 (2.90) 58.07 (1.08) 40.82 (1.27) 19.49 (0.18) 21.04 (0.47) 58.10 41.55
+IDAS 72.68 (0.20) 52.11 (0.60) 85.12 (0.14) 69.72 (0.29) 65.58 (0.24) 38.33 (0.45) 65.14 (0.32) 46.77 (0.85) 23.39 (0.36) 22.48 (0.45) 62.38 45.88
+SPILL 63.73 (0.67) 42.36 (1.14) 59.72 (1.05) 36.74 (0.92) 49.54 (1.41) 28.26 (0.94) 43.26 (0.85) 28.06 (1.28) 9.51 (0.44) 13.48 (0.20) 45.15 29.78
+LeBot (Ours) 80.26 (0.28) 65.45 (1.43) 88.97 (0.11) 77.70 (1.03) 70.15 (0.23) 40.00 (1.30) 69.60 (0.06) 54.52 (0.69) 27.89 (0.30) 25.82 (0.34) 67.37 52.70
+Bert
+Plain 50.72 (0.42) 29.07 (0.85) 72.90 (0.60) 49.48 (0.83) 62.24 (0.46) 26.78 (0.32) 49.12 (0.64) 30.49 (1.00) 10.77 (0.46) 12.90 (0.37) 49.15 29.75
+IDAS 70.30 (0.32) 49.85 (0.49) 87.97 (0.27) 71.57 (1.02) 68.90 (0.29) 33.08 (0.69) 65.24 (0.32) 49.06 (0.59) 14.10 (0.52) 15.22 (0.73) 61.30 43.76
+IDAS 72.73 (0.26) 53.45 (0.82) 88.70 (0.18) 74.99 (0.61) 69.52 (0.21) 33.19 (0.87) 66.41 (0.44) 50.04 (0.94) 14.89 (0.26) 16.50 (0.52) 62.45 45.63
+SPILL 65.67 (0.49) 43.44 (1.02) 86.91 (0.28) 71.39 (1.28) 69.50 (0.28) 32.49 (0.40) 62.67 (0.29) 43.64 (0.81) 12.30 (0.32) 13.86 (0.44) 59.41 40.96
+LeBoT (Ours) 76.86 (0.24) 60.95 (1.11) 92.90 (0.05) 84.47 (0.46) 72.28 (0.24) 38.69 (1.46) 73.16 (0.27) 61.02 (0.82) 19.25 (0.19) 18.77 (0.14) 66.89 52.78
+MiniLM
+Plain 79.08 (0.80) 61.27 (1.36) 89.46 (0.19) 73.73 (0.95) 67.41 (0.53) 31.40 (1.62) 70.37 (0.62) 54.04 (1.28) 10.84 (0.16) 13.89 (0.63) 63.43 46.87
+IDAS 83.01 (0.19) 65.96 (1.19) 92.82 (0.22) 80.05 (0.21) 70.46 (0.46) 35.55 (1.33) 75.67 (0.47) 59.52 (2.12) 22.27 (0.16) 27.22 (0.37) 68.85 53.66
+IDAS 84.88 (0.19) 72.77 (0.59) 93.41 (0.06) 85.52 (0.33) 70.88 (0.22) 36.46 (1.46) 75.93 (0.17) 57.75 (0.57) 21.25 (0.25) 25.67 (0.45) 69.27 55.63
+SPILL 81.92 (0.16) 66.22 (0.46) 90.78 (0.22) 77.76 (0.07) 70.48 (0.41) 37.95 (1.27) 72.59 (0.31) 58.05 (1.40) 15.54 (0.31) 17.22 (0.75) 66.26 51.44
+LeBoT (Ours) 85.44 (0.18) 73.12 (1.10) 94.45 (0.08) 87.58 (0.70) 72.18 (0.29) 39.40 (0.83) 76.74 (0.24) 66.84 (0.67) 21.77 (0.12) 25.22 (0.74) 70.12 58.43
+E5
+Plain 77.86 (0.35) 60.88 (0.89) 91.09 (0.17) 75.70 (0.36) 71.17 (0.28) 33.04 (0.74) 71.70 (0.84) 53.97 (2.00) 21.13 (0.44) 21.93 (0.60) 66.59 49.10
+IDAS 83.91 (0.24) 68.78 (1.11) 93.53 (0.26) 82.20 (1.22) 73.58 (0.37) 38.87 (1.04) 76.70 (0.89) 59.61 (3.25) 26.30 (0.17) 29.56 (0.64) 70.80 55.80
+IDAS 83.37 (0.14) 70.01 (0.59) 94.07 (0.09) 85.38 (0.35) 73.09 (0.45) 39.84 (1.24) 77.29 (0.15) 63.64 (0.80) 25.99 (0.09) 28.72 (1.08) 70.76 57.52
+SPILL 83.56 (0.47) 70.25 (1.56) 92.93 (0.08) 83.18 (0.78) 71.77 (0.35) 36.83 (1.10) 75.40 (0.51) 60.28 (1.81) 25.14 (0.36) 24.82 (0.86) 69.76 55.07
+ClusterLLM 84.16 (0.36) 70.13 (1.34) 92.92 (0.29) 80.48 (0.93) 74.46 (0.11) 37.22 (1.18) 74.39 (0.21) 56.08 (1.01) 22.23 (0.17) 22.22 (1.15) 69.63 53.23
+LeBoT (Ours) 85.34 (0.06) 73.86 (0.76) 93.63 (0.06) 85.65 (0.46) 72.51 (0.32) 40.53 (1.64) 76.16 (0.17) 62.75 (0.67) 26.70 (0.04) 33.70 (0.26) 70.87 59.30
+Table 3: Five-benchmark results (emerging data scenario).
+Averages over 5 runs (±SD).
+Gemma is used in LeBoT, SPILL and IDAS.
+Plain: direct embedding clustering.
+Bold numbers: best within the embedder.
+ClusterLLM cited directly from prior work.
+is with access to information of known $K$ during representation refinements.
+GCD scenario
+Table 4 shows the results for GCD.
+The goal is to examine how increasing the amount of data affects performance and to see the gap between our approach and semi-supervised learning.
+Our method achieves the best performance on all benchmarks in the fully unlabeled setting.
+It even outperforms baselines that know $10\%$ of the categories, and is comparable to other baselines that know $25\%$ of the categories (or up to $75\%$ in the case of IntentGPT).
+These strong results suggest that our approach can substantially reduce the need for human labeling.
+Bank 77 Clinc150 StackO Average
+KCR Method NMI Acc NMI Acc NMI Acc NMI Acc
+$\mathbf{0\%}$
+IntentGPT 81.42 64.22 94.35 83.20 - - - -
+LeBoT (Ours) 86.90 74.93 95.32 88.71 81.21 82.54 87.81 82.06
+$\mathbf{10\%}$
+LOOP 79.14 64.97 93.52 84.89 75.98 80.50 82.88 76.79
+GLEAN 82.23 67.99 95.21 88.71 79.67 82.40 85.70 79.70
+$\mathbf{25\%}$
+ALUP 84.06 74.61 94.84 88.40 76.58 82.20 85.16 81.74
+LOOP 83.37 71.40 94.38 86.58 79.10 82.20 85.62 80.06
+GLEAN 85.62 76.98 96.27 91.51 80.90 84.10 87.60 84.20
+$\mathbf{75\%}$ IntentGPT 85.94 77.21 96.06 88.76 - - - -
+Table 4: Three-benchmark results (GCD scenario).
+Averages over 5 runs.
+KCR: Known Categories Ratio.
+All-MiniLM-L6-v2 used as in LeBoT.
+Bold: best.
+Underline, second best.
+StackO: StackOverflow dataset.
+GLEAN and ALUP cited directly from prior work.
+LOOP cited directly from GLEAN.
+is with access to information of known $K$ during representation refinements.
+6.2 Further analysis
+Dimensionality of the bag-of-text representations
+We observe that the increase in dimensionality by $d^{*}$ is negligible: for most runs, the values fall between 0 and 20, and the maximum increase across all runs, datasets, and backbone embedders is only 44.
+Since we start with a relatively large initial dimension of $d=1,024$, this increase is very limited.
+This is expected, as each text is more likely to identify similar counterparts when the pool of candidate texts is large.
+We also tested other initial dimensions, and Figure 2 shows that the results remain robust across different initial settings.
+Figure 2: ACC and NMI across all test datasets (Emerging Data and GCD scenarios) for different $d$. $4,096$ means $\min\{4096,N\}$ since some datasets are smaller.
+Figure 3: Comparison of (a) varying neighbor $m$ and (b) iterative updates on ACC and NMI across all test datasets (Emerging Data and GCD). Backbone embedder: All-MiniLM-L6-v2.
+Effect of the Number of subset candidates m and Iterative Updates
+We use $m=30$ across all experiments, but this raises the question of whether different values of $m$ would affect performance.
+Figure 3(a) shows that small size of $m=10$ leads to lower results because the LLM has too few candidates to choose from.
+However, once $m\geq 20$, performance remains largely stable.
+To investigate the effect of iterative updates, Figure 3(b) shows that in subsequent iterations, the results substantially improve, especially after the first iteration, but no longer after the or third.
+Since most BoT vectors converge within 5 iterations, there are only minimal changes afterward: We find that most datasets have 90% convergence ratio by the 5th iteration (see Appendix G).
+Effect of Different LLMs
+Our main results were obtained with gemma-2-9b-it.
+For completeness, we also experimented with alternative LLMs.
+Importantly, because our objective is to evaluate whether the vectors encode LLM-derived selections rather than to benchmark LLM performance, these analyses are presented as ancillary robustness checks.
+Figure 4 shows the results obtained with three different LLMs.
+The results between Qwen and Gemma are similar.
+There is a slight drop with Llama.
+This finding is consistent with prior work Lin et al.
+(2025), as Llama produces more incorrect selections (see Appendix H for details).
+Figure 4: ACC and NMI across all test datasets (Emerging Data and GCD) for three LLMs. Backbone embedder: All-MiniLM-L6-v2
+Evaluation with different clustering algorithms
+Our approach does not require the number of clusters $K$ when building the representations.
+We therefore also evaluate it with HDBSCAN, which estimates the number of clusters automatically and is therefore closer to real-world use.
+Figure 5 shows that performance remains similar, indicating that our method is robust across different clustering algorithms.
+This confirms its practical applicability.
+Detailed implementation in Appendix I.
+Figure 5: ACC and NMI across all test datasets (Emerging Data and GCD) for HDBSCAN and K-means. Backbone embedder: All-MiniLM-L6-v2
+Increasing scalability with distillation
+[…]
+7 Conclusions
+We propose LeBoT (LLM-enabled Bag-of-Texts), an intuitive, domain-adaptive, label-free, and training-free method for short text clustering.
+Compared with other SOTA methods, LeBoT is less dependent on careful selection of the embedder and does not assume prior knowledge of clusters or labels, making it more aligned with real-world scenarios.
+LeBoT achieves better or similar results to state-of-the-art approaches, including methods that require labeled data.
+Experiments on diverse datasets shows our method is the most robust across different embedders, and experiments with smaller LLMs show that our method is robust in low-resource settings, while also scalable to large datasets.
+In the future, we plan to incorporate human feedback into our method pipeline.
+Bank 77 Clinc150
+Method NMI Acc NMI Acc
+K-means
+Plain 76.31 59.41 88.99 73.99
+LeBoT (Ours) 81.44 68.99 91.19 83.62
+HDBSCAN
+Plain 76.15 60.27 84.31 62.84
+LeBoT (Ours) 81.35 68.01 91.50 85.28
+Table 5: Clustering results on the entire large datasets using knowledge distillation. Averages over 5 runs for K-means. Plain: Directly embedding the test dataset with All-MiniLM-L6-v2.
+Acknowledgments
+This publication is part of the project LESSEN with project number NWA.1389.20.183 of the research program NWA ORC 2020/21 which is (partly) financed by the Dutch Research Council (NWO).
+Limitations
+Language.
+Like most of the prior work, we only focus on English utterance datasets.
+This is relevant because most of benchmark datasets are in English.
+Our method is completely data-driven and not dependent on labeled data and it is therefore applicable to other languages than English.
+LLM capability.
+Our method focuses on encoding preferences through Bag-of-Text vectors.
+Its effectiveness naturally depends on the LLM, and in highly specialized domains, some in-context learning may be required to improve the selection.
+Embedder capability.
+Although we construct a Bag-of-Text vector that reduces reliance on the existing embedder, for computational efficiency we still retrieve subset of candidate texts using the embedder.
+This will affect the quality of the constructed vectors if the embedder is very poor.
+This can affect the quality of the constructed vectors when the embedder is very weak.
+However, our main results show that this influence is relatively small compared with other methods.
+Appendix A Impact of our computational method on preference preservation
+In this section, we analyze the impact of dimension reduction and subset selection on the preservation of preference, assuming that the stated Assumption 1 hold.
+We do not discuss pooling here, because pooling is intended to handle cases where Assumption 1 is violated, and to account for the estimation involved in subset selection.
+Definition 2 (Preservation of Preference).
+We say that a function preserves preferences if $l_{x,x^{+}}>l_{x,x^{-}}$, where $l$ denotes the cosine similarity, $(x,x^{+})$ is a similar pair, and $(x,x^{-})$ is a disimilar pair.
+We need a generalized version of $g$ to consider the dimension reduction and subset selection.
+Definition 3 (Bag-of-Texts: generalization version of $g$).
+Let $g^{*}(x,\mathcal{T},\mathcal{L}(x))$ be a function that takes:
+  • a text $x$,
+  • a set of texts $D$,
+  • a subset of texts $\mathcal{T}\subseteq D$,
+  • a subset of texts $\mathcal{L}(x)\subseteq\mathcal{T}$, selected based on $x$
+and outputs a BoT representation $\mathbf{z}\in\mathbb{R}^{|\mathcal{T}|}$.
+In this representation, the entry corresponding to each text in $\mathcal{L}(x)$ that is considered similar to $x$ is assigned a value of 1, while all other entries are 0.
+The definition of the generalized function $g^{*}$ is to consider the reduced dimension of BoT (memory constraints) and subset selection (computation constraints). Note $g^{*}(x,D,D,D)$ = $g(x,D)$.
+Proposition 2 (Preservation of Preference on generalization function $g^{*}$).
+Let $D^{r}\subseteq D$ be such that for every cluster $S\in\mathcal{S}$, there exists at least one text $y\in D^{r}$ with $y\in S$.
+Let $M^{*}=\mathcal{L}(x)\subseteq\mathcal{T}$ such that, for every cluster $S\in\mathcal{S}$, there exists at least one text $o_{S}\in S$ that is included in $M^{*}$ for $x\in S$.
+$g^{*}(x,D,D^{r},M^{*})$ is a function that preserves the preference under Assumption 1.
+In the preposition 2, the dimensions constructed using $D^{r}$ ensure that each cluster has at least one dimension in which a value can be assigned.
+The subset $M^{*}$ guarantees that BoT vectors corresponding to texts within the same cluster share a 1 in at least one of these dimensions.
+As a result, the cosine similarity between vectors of texts from the same cluster is strictly positive due to the shared dimension(s).
+Furthermore, the more shared texts from the cluster that appear in $M^{*}$ for each $x$, the greater the overlap of 1s, and thus the higher the cosine similarity.
+This proposition 2 implies two key insights.
+First, instead of turning every text $x\in D$ into a dimension of the BoT vector, we may use only a subset, provided that the subset is sufficiently representative.
+Second, scanning the entire dataset for every $x$ can be replaced with scanning only such a subset, as long as it contains at least one shared element for each cluster.
+Remark 1: The proposition 2 is primarily intended for conceptual understanding.
+The shared text $o_{S}$ and subset of texts $M^{*}_{j}$ can also include the texts from remaining texts $D\setminus D^{r}$ not just those in $D^{r}$ since these remaining texts are initially constructed with $D^{r}$.
+This increases the number of shared texts available across clusters, providing more opportunities for matching and potentially improving similarity estimates.
+Remark 2: As mentioned in the section 3 and 4, in practice, we select a fairly large initial dimension $d$ and use an LLM to include missing representative texts to form $D^{r}$.
+For subset selection, for each $x_{j}$, we use pre-trained embeddings to select the top $m$ closest texts, denoted $M_{j}$, to approximate $M^{*}_{j}$.
+Appendix B Algorithm for Updating the Bag-of-Texts Vectors
+The BoT vector update largely involves a repetitive process similar to the original construction. See Algorithm 3 for details.
+Algorithm 3 Update Bag-of-Texts Vectors
+0:$D$, $\mathbf{X}$, $\mathbf{Z}$, candidates size $m$, max iterations $T$, LLM
+0: Bag-of-texts vectors $\mathbf{Z}\in\mathbb{R}^{N\times(d+d^{*})}$
+1: Initialize $\mathbf{Z}^{0}\leftarrow\mathbf{Z}$
+2:for $t=0$ to T1T-1 do
+3:  $\mathbf{U}^{t}\leftarrow\mathbf{X}\oplus\mathbf{Z}^{t}$
+4:  for each $x_{j}\in D$ do
+5:   if $t>0$ and $\cos(\mathbf{z}_{j}^{t-1},\mathbf{z}_{j}^{t})>0.99$ then
+6:    $\mathbf{z}_{j}^{t+1}\leftarrow\mathbf{z}_{j}^{t}$ {# Skip updating}
+7:   else
+8:    $M_{j}^{t}\leftarrow\{x^{t}_{j1},\dots,x^{t}_{jm}\}$ from $\mathbf{U}^{t}$ as the top-$m$ by cosine similarity to $x_{j}$
+9:    $O^{t}_{j}\leftarrow\text{LLM}(M_{j}^{t})\in\mathcal{P}(M_{j})\cup\{\emptyset\}$ {# $\mathcal{P}(.)$: power set}
+10:    if $O^{t}_{j}\neq\emptyset$ then
+11:     $\mathbf{z}^{t+1}_{j}\leftarrow\text{MeanPooling}\{\mathbf{z}^{t}_{k}:x_{k}\in O^{t}_{j}\}$
+12:     $\mathbf{z}^{t+1}_{j}\leftarrow\frac{\mathbf{z}^{t+1}_{j}}{\|\mathbf{z}^{t+1}_{j}\|_{2}}$
+13:    else
+14:     $\mathbf{z}_{j}^{t+1}\leftarrow\mathbf{z}_{j}^{t}$
+15:    end if
+16:   end if
+17:  end for
+18:  $\mathbf{Z}^{t+1}\leftarrow\{\mathbf{z}_{j}^{t+1}\}_{x_{j}\in D}$
+19:end for
+20:return $\mathbf{Z}^{T}\in\mathbb{R}^{N\times(d+d^{*})}$
+Appendix C Comparison of Data Splits in GCD scenario: Our Method vs. Prior Work
+Table 6 shows the train–validation–test splits used in ALUP (Liang et al., 2024b) and GLEAN Zou et al.
+(2025b) for the three benchmark datasets.
+In their setting, a specified ratio of all categories is randomly selected as known categories, referred to as the known category ratio (KCR).
+For each known category, 10% of the data is chosen to form the labeled dataset [^1] IntentGPT (Rodriguez et al., 2024) mention they use fewer than 10% for each known category, but they do not specify the exact percentage., while the remaining samples constitute the unlabeled dataset.
+Together, these two subsets form the training dataset.
+In our experiments, we do not use the validation set and use only 30% of the training dataset, restricted to the unlabeled setting with no access to labeled data.
+Evaluation is conducted on the same test split.
+They train an embedder using the training split and optimize it based on the labeled validation set, with the test set used solely for evaluation.
+In contrast, our method performs test-time embedding construction, so it requires to access test data, similar to IDAS De Raedt et al.
+(2023) and IntentGPT Rodriguez et al.
+(2024), using all available text data ($D=D^{test}\cup D^{u}$) to construct embeddings.
+Since our clustering is unsupervised, no label information is used.
+In practice, clustering methods require access to the text to be clustered in order to construct representations, and our approach naturally uses all available texts.
+# clusters # train # valid # test # total
+Bank77 77 9,003 1,000 3,080 13,083
+Clinc150 150 18,000 2,250 2,250 22,500
+StackO 20 18,000 1,000 1,000 20,500
+Table 6: Train–Validation–Test splits from previous studies (An et al., 2024; Liang et al., 2024b; Zou et al., 2025b). ‘StackO’ is the StackOverflow dataset
+Appendix D Selection of $d$
+gemma-2-9b-it (as LLM) and All-MiniLM-L6-v2 (as embedder) are used in the hyperparameter search.
+Table 7 shows the two datasets that we used for selecting the dimensionality of the BoT vector $d$.
+Figure 6 shows that they all give similar performance.
+We selected value $d=1024$ and used this value across all datasets and experiments.
+[…]
+ArxivS2S 93 3,674
+Reddit 50 3,217
+Table 7: External datasets used for selecting the dimensionality hyperparameter $d$
+Figure 6: The average ACC and NMI across two datasets for different $d$. Results for K-means are averaged over 5 runs. 40964096 here means $\min\{4096,N\}$ because the dataset size can be smaller.
+Appendix E Prompt
+Table 8 shows the prompts used in our experiment. It consists of three parts: task instruction, answer format, and task.
 CLINC150, Bank77, Mtop, and Massive
@@
 If the dataset is Stackoverflow, we use ‘question’ instead of ‘utterance’ in the Answer format.
-5.3 Hyperparameter setting
-The main hyperparameter in our approach is the initial dimension (number of representatives) of the sparse vector $d$, which we set to 2,048 based on validation results on two external datasets: ArxivS2S and Reddit (Muennighoff et al., 2023).
-We use external datasets to prevent overfitting on a specific dataset.
-See details in Appendix B.
-Qwen3-8B (as LLM) and All-MiniLM-L6-v2 (as embedder) are used in the hyperparameter search.
-5.4 Evaluation
-The emerging and GCD scenarios are evaluated under two conditions: with HDBSCAN (McInnes et al., 2017) and K-means (Lloyd, 1982).
+Appendix F Effect of Different Embedders
+Because we use a subset instead of scanning the entire dataset for each text, the quality of the texts selected by the embedder influences the construction of our BoT vector.
+We report two ratios of the first iteration, averaged across all texts.
+Embedder Selection Ratio:
+$\frac{\#\{\text{True same-cluster in top $m$ as the text}\}}{m}$
+In our experiments, we set $m=30$.
+Note that Embedder here does not refer to the pre-training embeddings alone; it denotes the concatenation of the pre-training embeddings and the BoT vectors.
+We focus on the first iteration, i.e., $\mathbf{U}^{1}$, since all BoT vectors have not yet converged and they are largely evenly spaced and uninformative in the initial state.
+As a result, the selection relies mostly on the pre-training embeddings.
+LLM Selection Ratio (LLM):
+[…]
+Table 9 shows that E5 produces the best pool on average. We also observe that the LLM ratio remains relatively stable, even when the embedder ratio fluctuates substantially.
+Bank 77 Clinc150 Mtop Massive GoEmo Average
+Emb LLM Emb LLM Emb LLM Emb LLM Emb LLM Emb LLM
+TF-IDF 43.26 68.02 49.71 82.60 62.93 78.55 45.28 70.21 30.85 36.32 46.41 67.14
+Bert 38.08 63.41 56.12 86.49 69.72 82.48 50.48 73.61 20.02 25.90 46.88 66.38
+MiniLM 62.03 75.55 69.76 89.09 70.88 81.27 61.58 76.14 22.33 27.79 57.32 69.97
+E5 58.64 73.25 67.80 88.48 72.00 81.53 61.66 76.33 27.97 32.02 57.61 70.32
+Table 9: Selection Accuracy By Embedders at 1st iteration ($\%$)(emerging data scenario). Bold numbers: best. LLM is Gemma.
+Appendix G Percentage of Texts Converged at Iteration $t$
+Figure 7 shows ratio of text convergence. It shows most datasets have 90% convergence ratio by the 5th iteration, and by the 7th iteration, all datasets achieve a 95% convergence ratio.
+Figure 7: Ratio of texts converged at iteration $t$. All-MiniLM-L6-v2 (as embedder) and Qwen3-8B (as LLM) were used.
+Bank 77 Clinc150 Mtop Massive GoEmo Bank 77GCD Clinc150GCD StackOGCD Average
+NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc
+Gemma 85.44 73.12 94.45 87.58 72.18 39.40 76.74 66.84 21.77 25.22 86.90 74.93 95.32 88.71 81.21 82.54 76.75 67.29
+Qwen 85.01 71.63 93.79 87.12 73.29 37.45 74.24 61.47 21.61 25.83 86.53 74.31 95.61 89.41 82.15 83.42 76.53 66.33
+Llama 81.75 66.18 91.66 83.12 70.12 35.79 73.26 60.40 18.91 20.81 83.16 68.16 94.00 85.66 77.42 80.10 73.78 62.53
+Table 10: Further analysis of results using different LLMs.
+Appendix H Results by Different LLMs
+We report the detailed results of all datasets by different LLMs in Table 10.
+Qwen and Gemma consistently outperform LLaMA.
+We also report the LLM selection accuracy (see Appendix F for the definition).
+Table 11 shows that Gemma and Qwen have similar capabilities in identifying similar texts, outperforming LLaMA.
+This finding aligns with prior work Lin et al.
+(2025).
+Bank 77 Clinc150 Mtop Massive GoEmo Average
+Emb LLM Emb LLM Emb LLM Emb LLM Emb LLM Emb LLM
+Gemma 62.03 75.55 69.76 89.09 70.88 81.27 61.58 76.14 22.33 27.79 57.32 69.97
+Qwen 61.08 75.32 70.16 90.14 72.44 84.19 60.88 76.62 22.27 28.91 57.36 71.04
+Llama 56.15 64.57 64.05 78.95 67.99 75.16 58.17 67.74 19.96 23.43 53.26 61.97
+Table 11: Selection Accuracy By LLMs at 1st iteration($\%$)(emerging data scenario). Bold numbers: best. All-MiniLM-L6-v2 used as backbone embedder.
+Appendix I HDBSCAN Clustering
 To reflect real-world settings with unknown $K$, we apply HDBSCAN (McInnes et al., 2017) to partition $D^{test}$ into $\hat{K}$ clusters.
@@
 The noise points are assigned to the nearest cluster centroid.
-We perform an incremental grid search over min_cluster_size values of $10,15,\ldots,50$, stopping early if the silhouette score does not improve (Rousseeuw, 1987; Gung et al., 2023).
-K-means clustering is the dominant evaluation practice because many studies assume a known cluster number $K$ and incorporate it in their method pipelines Zhang et al.
-(2023); De Raedt et al.
-(2023); Liang et al.
-(2024b); Feng et al.
-(2024); Zou et al.
-(2025b).
-Following their convention, the number of clusters for K-means is set to match the number of ground-truth cluster number $K$.
-To reduce cost, we follow (Zhang et al., 2023) and perform a single pipeline run per dataset using a fixed seed to obtain a sparse vector for each text.
-Most steps are deterministic.
-The main source of variability comes from splitting candidate examples across multiple LLM calls (e.g., three calls with 10 examples instead of one with 30), done only to save time and GPU memory.
-LLM outputs are generated via greedy decoding, without sampling.
-For evaluation, K-means is run over 5 different seeds and averaged.
-HDBSCAN, which is near-deterministic, is run once.
-After we derive the clusters, we apply standard clustering metrics for evaluation.
-These metrics include normalized mutual information (NMI) and clustering accuracy (Acc) (Rand, 1971; Meilă, 2007; Huang et al., 2014; Gung et al., 2023).
-5.5 Baselines
-Emerging Data Scenario
-We compare against SCCL (Zhang et al., 2021), ClusterLLM, (Zhang et al., 2023), IDAS De Raedt et al.
-(2023), LLMedgeRefine (Feng et al., 2024), and SPILL (Lin et al., 2025).
-Table 3 shows that our method makes the fewest assumptions and has the lowest requirements.
-SPILL is the only method under the same assumptions.
-However, we still include ClusterLLM, IDAS, and LLMedgeRefine, which uses additional information – the known $K$ – in its development pipeline, to provide a more comprehensive comparison.
-GCD Scenario
-We compare against IDAS De Raedt et al.
-(2023), Loop (An et al., 2024), ALUP Liang et al.
-(2024b), IntentGPT Rodriguez et al.
-(2024) and Glean Zou et al.
-(2025b).
-Note that IntentGPT and IDAS can also be implemented only using unlabeled data.
-Table 3 shows that our method again makes the fewest assumptions and has the lowest requirements.
-Emerging data FT known $K$ Labeled data
-TWIST (Ours) -
-SCCL (Zhang et al., 2021) $\checkmark$ -
-ClusterLLM (Zhang et al., 2023) $\checkmark$ $\checkmark$ -
-LLMEdgeRefine (Feng et al., 2024) $\checkmark$ -
-IDAS (De Raedt et al., 2023) $\checkmark$ -
-SPILL (Lin et al., 2025) -
-GCD
-TWIST (Ours)
-Loop (An et al., 2024) $\checkmark$ $\checkmark$ $\checkmark$
-ALUP (Liang et al., 2024b) $\checkmark$ $\checkmark$ $\checkmark$
-IntentGPT (Rodriguez et al., 2024)
-Glean (Zou et al., 2025b) $\checkmark$ $\checkmark$ $\checkmark$
-Table 3: Baselines settings. FT: Finetuning. known $K$: Assuming known $K$ during the method pipeline development. Labeled data: Labeled data is required.
-6 Results
-6.1 Main results
-Emerging data scenario
-Table 4 shows the HDBSCAN results.
-Our approach substantially improves performance compared to directly using the embedder.
-[^1] The estimated number of clusters, $\hat{K}$, is provided in Appendix D.
-Moreover, the results from BERT (non-contrastive) also improve with our approach.
-For the K-means results, Table 5 shows that our approach generally outperforms IDAS and ClusterLLM on both metrics across different embedders, and achieves comparable results to LLMEdgeRefine.
-This indicates that, even without assuming prior knowledge of $K$ during method development – as the compared methods do – our approach remains effective and better reflects real-world practice.
-While SPILL achieves slightly higher NMI, our method attains substantially higher ACC.
-Lastly, we found that the HDBSCAN results overall are similar to those of K-means, indicating that our method remains robust even without knowing $K$.
+We perform an incremental grid search over min_cluster_size values of $10,15,\ldots,50$, stopping early if the silhouette score does not improve (Rousseeuw, 1987).
+Using the silhouette score for hyperparameter search follows prior work (Gung et al., 2023).
+Tables 12 and 13 shows the detailed evaluation result with HDBSCAN clustering.
 Bank 77 Clinc150 Mtop Massive GoEmo Average
 NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc
+TF-IDF
+Plain 53.06 33.34 56.55 32.64 39.79 22.83 41.26 26.09 11.56 16.10 40.44 26.20
+LeBoT (Ours) 79.92 65.23 89.29 79.58 71.14 46.62 69.81 56.46 30.93 24.47 68.22 54.47
 Bert
 Plain 0.84 1.69 73.03 50.87 63.95 40.97 2.86 8.18 1.35 12.05 28.41 22.75
-TWIST (Ours) 76.33 60.23 93.23 85.91 73.86 47.47 72.02 62.32 21.01 14.19 67.29 54.02
+LeBoT (Ours) 76.12 57.40 93.07 85.62 75.94 51.64 73.71 63.97 20.78 15.68 67.92 54.86
 MiniLM
 Plain 78.78 60.39 86.79 67.71 67.30 55.91 62.55 48.15 11.31 18.62 61.35 50.16
-TWIST (Ours) 85.61 75.29 93.86 86.44 74.52 50.36 76.67 67.51 21.63 23.27 70.46 60.57
+LeBoT (Ours) 85.13 71.85 94.52 87.58 76.83 59.51 76.88 67.78 23.18 21.92 71.31 61.73
 E5
 Plain 78.68 62.18 87.57 65.62 63.45 53.79 1.15 7.74 15.43 17.70 49.26 41.41
-TWIST (Ours) 84.86 72.24 93.41 86.89 72.53 44.32 75.85 65.12 27.77 29.03 70.88 59.52
-Instructor
-Plain 82.50 67.34 92.29 79.27 67.63 54.31 67.99 53.54 0.25 11.48 62.13 53.19
-TWIST (Ours) 85.63 73.77 93.64 84.62 74.06 46.81 76.52 64.07 31.64 33.65 72.30 60.58
-Table 4: Five-benchmark results (Emerging Data Scenario, HDBSCAN). Gemma is used in TWIST. Plain: direct embedding clustering. Bold numbers: best within the embedder.
-Bank 77 Clinc150 Mtop Massive GoEmo Average
-NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc
-Bert
-Plain 50.72 (0.42) 29.07 (0.85) 72.90 (0.60) 49.48 (0.83) 62.24 (0.46) 26.78 (0.32) 49.12 (0.64) 30.49 (1.00) 10.77 (0.46) 12.90 (0.37) 49.15 29.75
-TWIST (Ours) 76.23 (0.14) 60.49 (0.72) 93.25 (0.07) 86.43 (0.65) 71.79 (0.20) 36.41 (0.20) 72.00 (0.13) 60.56 (0.61) 18.66 (0.10) 18.76 (0.29) 66.38 52.53
-MiniLM
-Plain 79.08 (0.80) 61.27 (1.36) 89.46 (0.19) 73.73 (0.95) 67.41 (0.53) 31.40 (1.62) 70.37 (0.62) 54.04 (1.28) 10.84 (0.16) 13.89 (0.63) 63.43 46.87
-TWIST (Ours) 85.57 (0.06) 74.07 (0.71) 93.65 (0.07) 86.35 (0.34) 71.92 (0.22) 37.83 (0.72) 76.18 (0.26) 66.15 (0.58) 20.69 (0.09) 24.44 (0.37) 69.60 57.77
-E5
-Plain 77.86 (0.35) 60.88 (0.89) 91.09 (0.17) 75.70 (0.36) 71.17 (0.28) 33.04 (0.74) 71.70 (0.84) 53.97 (2.00) 21.13 (0.44) 21.93 (0.60) 66.59 49.10
-SCCL 77.34 (0.62) 63.60 (1.37) 91.89 (0.49) 77.96 (1.78) 70.42 ((0.34) 33.82 ((1.07) 71.57 (0.89) 54.48 (1.80) 20.05 (0.63) 22.03 (0.69) 66.25 50.38
-ClusterLLM 84.16 (0.36) 70.13 (1.34) 92.92 (0.29) 80.48 (0.93) 74.46 (0.11) 37.22 (1.18) 74.39 (0.21) 56.08 (1.01) 22.23 (0.17) 22.22 (1.15) 69.63 53.23
-SPILL 83.56 (0.47) 70.25 (1.56) 92.93 (0.08) 83.18 (0.78) 71.77 (0.35) 36.83 (1.10) 75.40 (0.51) 60.28 (1.81) 25.14 (0.36) 24.82 (0.86) 69.76 55.07
-TWIST (Ours) 85.09 (0.19) 74.23 (0.88) 93.31 (0.04) 85.30 (0.58) 71.57 (0.18) 37.32 (1.67) 75.72 (0.22) 62.93 (1.28) 26.58 (0.23) 33.20 (0.54) 70.45 58.60
-Instructor
-Plain 82.42 (0.68) 65.49 (2.41) 93.27 (0.32) 81.68 (1.31) 71.63 (0.39) 33.14 (0.52) 74.02 (0.54) 54.62 (2.06) 29.18 (0.76) 32.51 (1.51) 70.10 53.49
-SCCL 81.77 (1.36) 65.48 (1.36) 92.94 (0.44) 80.85 (0.74) 73.52 (0.38) 34.28 (0.58) 73.90 ((0.36) 54.10 ((1.05) 30.54 (0.68) 34.33 (0.86) 70.53 53.81
-ClusterLLM 85.15 (0.41) 71.20 (1.59) 94.00 (0.21) 83.80 (0.41) 73.83 (0.79) 35.04 (0.97) 77.64 (0.21) 60.69 (0.96) 23.89 (0.68) 26.75 (1.76) 70.90 55.50
-LLMEdgeRefine - (0.00) - (0.00) 94.86 (0.00) 86.77 (0.00) 72.92 (0.00) 46.00 (0.00) 76.66 (0.00) 63.42 (0.00) 29.74 (0.00) 34.76 (0.00) - -
-IDAS - (0.00) - (0.00) 92.35 (0.00) 81.36 (0.00) 72.31 (0.00) 37.30 (0.00) 75.74 (0.00) 63.01 (0.00) 25.57 (0.00) 30.61 (0.00) - -
-SPILL 85.01 (0.29) 71.05 (0.83) 93.77 (0.35) 85.14 (1.04) 72.65 (0.32) 37.11 (1.62) 77.62 (0.46) 62.42 (2.06) 32.78 (0.52) 33.21 (2.33) 72.37 57.79
-TWIST (Ours) 85.46 (0.18) 74.66 (1.01) 93.92 (0.09) 85.62 (0.71) 72.34 (0.28) 38.40 (0.81) 76.43 (0.03) 63.17 (0.25) 30.62 (0.24) 37.04 (1.55) 71.75 59.78
-Table 5: Five-benchmark results (Emerging Data Scenario, K-means).
-Averages over 5 runs (±SD).
-Gemma is used in TWIST.
-Plain: direct embedding clustering.
-Bold numbers: best within the embedder.
-SCCL/ClusterLLM/LLMEdgeRefine cited directly from prior work.
-SCCL results are cited from ClusterLLM; IDAS results are cited from LLMedgeRefine.
-GCD scenario
-Table 6 shows the HDBSCAN results for GCD.
-Our method consistently improves the performance of the existing embedder.
-Table 7 shows the K-means results for GCD.
-Our approach is on average the best in the unlabeled setting, without any known categories.
-It even outperforms baselines that know $10\%$ of the categories, and is comparable to other baselines that know $25\%$ of the categories (or up to $75\%$ in the case of IntentGPT).
-These strong results, together with the fewer assumptions required by our method, highlight its effectiveness and efficiency.
-Notably, BANK77 and CLINC150 perform similarly with both clustering algorithms, while Stackoveflow shows a noticeable drop.
+LeBoT (Ours) 85.03 70.65 93.82 85.80 75.38 53.17 76.33 62.59 28.07 29.59 71.73 60.36
+Table 12: Five-benchmark results (emerging data scenario, HDBSCAN). Gemma is used in LeBoT. Plain: direct embedding clustering. Bold numbers: best within the embedder.
 Bank 77 Clinc150 StackO Average
-KCR Method NMI Acc NMI Acc NMI Acc NMI Acc
+Method NMI Acc NMI Acc NMI Acc NMI Acc
 Plain 78.78 60.39 83.02 51.82 66.26 57.20 76.02 56.47
-$\mathbf{0\%}$ TWIST 87.43 75.16 95.19 89.07 78.54 71.70 87.05 78.64
-Table 6: Three-benchmark results (GCD scenario, HDBSCAN) Gemma is used in TWIST.
-Bold numbers: best.
-KCR ratio of known category.
-All-MiniLM-L6-v2 used as backbone embedder.
-Plain: direct embedding clustering.
-StackO: StackOverflow dataset.
-Bank 77 Clinc150 StackO Average
-KCR Method NMI Acc NMI Acc NMI Acc NMI Acc
-- Plain 79.08 61.27 89.46 73.73 72.49 72.68 80.34 69.08
-$\mathbf{0\%}$
-IDAS 82.84 67.43 93.82 85.48 81.26 83.82 85.97 78.91
-IntentGPT 81.42 64.22 94.35 83.20 - - - -
-TWIST (Ours) 87.67 77.10 95.21 89.07 81.69 82.56 88.19 82.91
-$\mathbf{10\%}$
-Loop 79.14 64.97 93.52 84.89 75.98 80.50 82.88 76.79
-GLEAN 82.23 67.99 95.21 88.71 79.67 82.40 85.70 79.70
-$\mathbf{25\%}$
-ALUP 84.06 74.61 94.84 88.40 76.58 82.20 85.16 81.74
-Loop 83.37 71.40 94.38 86.58 79.10 82.20 85.62 80.06
-GLEAN 85.62 76.98 96.27 91.51 80.90 84.10 87.60 84.20
-$\mathbf{75\%}$ IntentGPT 85.94 77.21 96.06 88.76 - - - -
-Table 7: Three-benchmark results (GCD scenario, K-means).
-Averages over 5 runs with Gemma.
-KCR: known categories ratio.
-All-MiniLM-L6-v2 used as backbone embedder in TWIST.
-Loop results are cited from GLEAN.
-Plain: direct embedding clustering.
-6.2 Further analysis
-Effect of Different LLMs
-Our main results were obtained with gemma-2-9b-it.
-To investigate whether using a different LLM affects performance, Figure 2 shows the results obtained with all-MiniLM-L6-v2 as the embedder and three different LLMs (see Appendix E for details).
-Overall, the differences between Qwen3-8B and gemma-2-9b-it are similar but an obvious drop with Llama3.1-8B-Instruct.
-Because Qwen3-8B achieves comparable results, we use it along with all-MiniLM-L6-v2 for remaining analyses to reduce computational cost.
-Figure 2: ACC and NMI across all test datasets (Emerging Data and GCD) for three LLMs. Results for K-means are averaged over 5 runs, HDBSCAN from a single run.
-Effect of the Number of Closest Neighbors m and Iterative Updates
-We use $m=30$ across all the experiment, but this raises the question of whether different values of $m$ would affect performance.
-Figure 3(a) shows that using too few neighbors leads to lower results.
-To investigate the effect of iterative updates, Figure 3(b) shows that in subsequent iterations, the results substantially improve, especially after the first iteration.
-Since most sparse vectors converge within 5 iterations, there are only minimal changes afterward: We find that most datasets have 90% convergence ratio by the 5th iteration (see Appendix F).
-Figure 3: Comparison of (a) varying negihbor $m$ and (b) iterative updates on ACC and NMI across all test datasets (Emerging Data and GCD).
-A scalable implementation of our method
-Our representation relies on LLM guidance for each text, which can be computationally expensive when the dataset is extremely large.
-To assess scalability, we additionally explore clustering over the full dataset.
-The goal is to implement our approach on a subset of the texts and then use a model to generalize to the remaining texts (a form of knowledge distillation), without requiring further LLM computation.
-Specifically, we train a lightweight MLP ($\sim$10M parameters) to map pre-trained embeddings to sparse vectors.
-See Appendix G for experiment details.
-Note that we do not use any labeled data.
-Table 8 shows that our approach still improves the original embeddings by a large margin, which proves the scalability of our method, offering substantial savings in time and computational resources.
-7 Conclusions
-We propose TWIST (Short Text Clustering through Iterative Sparse Vector Updating with LLMs), an intuitive, domain-adaptive, label-free, and training-free method for short text clustering.
-Compared with other SOTA methods, TWIST is built under minimal assumptions and requirements.
-TWIST achieves better or similar results to state-of-the-art approaches without assuming prior knowledge of clusters or labels, making it more aligned with real-world scenarios.
-Experiments on diverse datasets and smaller LLMs show that our method is robust in low-resource settings, and also scalable to large datasets.
-In the future, we plan to expand our approach to clustering in specialized domains and explore ways to incorporate human feedback into our method pipeline.
-Bank 77 Clinc150 StackOverf Average
-Method NMI Acc NMI Acc NMI Acc NMI Acc
-Kmeans
-Plain 76.31 59.41 88.99 73.99 68.42 65.93 77.91 66.44
-TWIST (Ours) 83.80 73.58 92.78 87.27 72.98 76.90 83.19 79.25
-HDBSCAN
-Plain 76.15 60.27 84.31 62.84 63.01 61.80 74.49 61.64
-TWIST (Ours) 83.39 71.79 92.55 86.04 71.08 69.58 82.34 75.80
-Table 8: Clustering results on the entire dataset using knowledge distillation. Averages over 5 runs for K-means. Plain: Directly embedding the test dataset with All-MiniLM-L6-v2.
-Acknowledgments
-This publication is part of the project LESSEN with project number NWA.1389.20.183 of the research program NWA ORC 2020/21 which is (partly) financed by the Dutch Research Council (NWO).
-Limitations
-Language.
-Like most of the prior work, we only focus on English utterance datasets.
-This is relevant because most of benchmark datasets are in English.
-Our method is completely data-driven and not dependent on labeled data and it is therefore applicable to other languages than English.
-LLM capability. Our method relies on the abilities of LLMs. Although we empirically show that small LLMs can achieve good results, this is not guaranteed when the dataset comes from a rare domain.
-Embedder capability.
-Although we empirically show that our approach generally improves performance compared to directly using the embedder for clustering, if the embeddings provide a poor initial result, the improvement might be limited.
-Appendix A Comparison of Data Splits in GCD scenario: Our Method vs. Prior Work
-Table 9 shows the train–validation–test splits used in Loop (An et al., 2024), ALUP (Liang et al., 2024b) and GLEAN Zou et al.
-(2025b) for the three benchmark datasets.
-In their setting, a specified ratio of all categories is randomly selected as known categories, referred to as the known category ratio (KCR).
-For each known category, 10% of the data is chosen to form the labeled dataset [^2] IntentGPT (Rodriguez et al., 2024) mention they use fewer than 10% for each known category, but they do not specify the exact percentage., while the remaining samples constitute the unlabeled dataset.
-Together, these two subsets form the training dataset.
-In our experiments, we do not use the validation set and use only 30% of the training dataset, restricted to the unlabeled setting with no access to labeled data.
-Evaluation is conducted on the same test split.
-They train an embedder using the training split and optimizes it based on the labeled validation set, with the test set used solely for evaluation.
-In contrast, our method performs test-time embedding construction, so it requires to access test data, similar to IDAS De Raedt et al.
-(2023) and IntentGPT Rodriguez et al.
-(2024), using all available text data ($D=D^{test}\cup D^{u}$) to construct embeddings.
-Since our clustering is unsupervised, no label information is used.
-In practice, clustering methods require access to the text to be clustered in order to construct representations, and our approach naturally uses all available texts.
-# clusters # train # valid # test # total
-Bank77 77 9,003 1,000 3,080 13,083
-Clinc150 150 18,000 2,250 2,250 22,500
-StackO 20 18,000 1,000 1,000 20,500
-Table 9: Train–Validation–Test splits from previous studies (An et al., 2024; Liang et al., 2024b; Zou et al., 2025b). ‘StackO’ is the StackOverflow dataset
-Appendix B Selection of $d$
-Table 10 shows the two datasets that we used for selecting the dimensionality of the sparse vector $d$.
-Figure 4 shows that $1024$, $2048$, and $\min\{4096,N\}$ give similar performance for both K-means and HDBSCAN.
-We selected value $d=2048$ and used this value across all datasets and experiments.
-[…]
-ArxivS2S 93 3,674
-Reddit 50 3,217
-Table 10: External datasets used for selecting the dimensionality hyperparameter $d$
-Figure 4: The average ACC and NMI across two datasets for different $d$.
-Results for K-means are averaged over 5 runs, while HDBSCAN is shown from a single run.
-$4096$ here means $\min\{4096,N\}$ because the dataset size can be smaller.
-Appendix C Prompts for Instructor embedder
-Our method is built on top of existing embedders.
-For instructor-large, a prompt can be added to generate task-specific embeddings.
-We follow previous research (Su et al., 2022; Liang and Liao, 2023; Lin et al., 2025) and use mostly the same prompts.
-The prompts we used for the different datasets are listed in Table 11.
-Instructor:
-Bank77: Represent the bank purpose for retrieval:
-Mtop and Stackoverflow: Represent the sentence for retrieval:
-Clinc150 and Massive: Represent the sentence for retrieving the purpose:
-GoEmo: Represent an emotion sentence for retrieval:
-Table 11: The prompts used in Instructor.
-Appendix D Estimated Number of Clusters $\hat{K}$ by HDBSCAN
+LeBoT 86.75 74.64 95.24 88.36 81.43 83.50 87.81 82.17
+Table 13: Three-benchmark results (GCD scenario, HDBSCAN) Gemma is used in LeBoT. Bold numbers: best. All-MiniLM-L6-v2 used as backbone embedder. Plain: direct embedding clustering.
 Table 14 and Table 15 show the estimated number of clusters, $\hat{K}$, for the two scenarios.
 The results indicate that the refined embeddings produced by our method are generally closer to the ground truth than the pre-trained (Plain) embeddings.
-An exception is the Bank77 dataset.
+Except for the Bank77 and GoEmo datasets.
 Bank 77 Clinc150 Mtop Massive GoEmo
@@
 Plain 2 111 62 2 2
-TWIST 73 154 80 57 63
+LeBoT 69 144 77 61 51
 MiniLM
 Plain 78 139 17 27 11
-TWIST 68 150 62 56 44
+LeBoT 66 145 67 56 53
 E5
 Plain 73 134 9 2 3
-TWIST 70 142 82 57 58
-Instructor
-Plain 71 144 18 29 2
-TWIST 84 148 99 58 57
-True $K$ 77 150 102 59 27
-Table 12: Emerging scenario: Estimated Number of Clusters $\hat{K}$ by HDBSCAN. gemma-2-9b-it used in TWIST.
+LeBoT 68 147 64 63 60
+Table 14: Emerging scenario: Estimated Number of Clusters $\hat{K}$ by HDBSCAN. gemma-2-9b-it used in LeBoT.
 Bank 77 Clinc150 Stackoverflow
 Plain 78 82 13
-TWIST 84 142 30
+LeBoT 80 139 19
 True $K$ 77 150 20
-Table 13: GCD scenario: Estimated Number of Clusters $\hat{K}$ by HDBSCAN. All-MiniLM-L6-v2 used as backbone embedder. gemma-2-9b-it was used in TWIST.
-Bank 77 Clinc150 Mtop Massive GoEmo Bank 77GCD Clinc150GCD StackOGCD Average
-NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc
-Kmeans
-Gemma 85.57 74.07 93.65 86.35 71.92 37.83 76.18 66.15 20.69 24.44 87.67 77.10 95.21 89.07 81.69 82.56 76.57 67.20
-Qwen 85.13 73.18 93.59 86.96 72.23 39.03 74.55 63.64 20.12 22.80 87.04 75.78 94.80 88.21 80.69 82.32 76.02 66.49
-Llama 82.46 69.81 91.03 82.21 70.08 36.21 72.47 60.47 17.60 19.67 84.34 70.91 93.83 86.27 75.69 78.68 73.44 63.03
-HCBSCAN
-Gemma 85.61 75.29 93.86 86.44 74.52 50.36 76.67 67.51 21.63 23.27 87.43 75.16 95.19 89.07 78.54 71.70 76.68 67.35
-Qwen 85.04 73.90 93.56 86.42 74.42 50.11 74.39 62.36 22.10 19.81 86.81 75.29 94.78 88.18 78.07 71.50 76.15 65.95
-Llama 82.45 68.70 91.00 78.64 72.80 51.12 73.19 63.84 18.53 16.84 84.16 69.35 93.84 84.18 74.78 72.20 73.84 63.11
-Table 14: Further analysis of results using different LLMs.
-Appendix E Results by Different LLMs
-We report the detailed results of all datasets by different LLMs in Table 14.
-Qwen and Gemma consistently outperform LLaMA.
-In addition, the average differences between K-means and HDBSCAN are minimal across both metrics for each LLM.
-Appendix F Percentage of texts converged at iteration $t$
-Figure 5 shows ratio of text convergence. It shows most datasets have 90% convergence ratio by the 5th iteration, and by the 7th iteration, all datasets achieve a 95% convergence ratio.
-Figure 5: Ratio of texts converged at iteration $t$. All-MiniLM-L6-v2 (as embedder) and Qwen3-8B (as LLM) were used.
-Appendix G Details of the full dataset clustering
+Table 15: GCD scenario: Estimated Number of Clusters $\hat{K}$ by HDBSCAN. All-MiniLM-L6-v2 used as backbone embedder. gemma-2-9b-it was used in LeBoT.
+Appendix J Details of the full dataset clustering
 Datasets Used with LLM Guidance
 Table 16 shows the number of examples we use with the LLM.
-The number of examples with LLM guidance is the same as in Table 1, i.e.
-$|D^{test}|+|D^{u}|$, so we do not need to re-run all experiments.
-The examples account for 29.6% to 44.2% of datasets.
-Examples with LLM guidance are used to train the MLP, since their sparse vectors have been derived.
-The pre-trained embeddings from all-MiniLM-L6-v2 serve as input, and the LLM-generated sparse vectors serve as target output for supervised learning with an MLP.
+We re-use the representations derived in the emerging data setting, so the number of examples with LLM guidance remains the same as in Table 1, i.e., $|D^{test}|$, and there is no need to re-run all experiments.
+The examples account for 20.6% to 23.5% of datasets.
+Examples with LLM guidance are used to train the MLP, since their BoT vectors have been derived.
+The pre-trained embeddings from all-MiniLM-L6-v2 serve as input, and the LLM-generated BoT vectors serve as target output for supervised learning with an MLP.
 20% of the examples are used as a validation set to tune the hyperparameters.
@@
 Clustering of the entire dataset with MLP output
-We first encode all examples into pre-trained embeddings using all-MiniLM-L6-v2, then use the trained MLP to derive sparse versions of these embeddings.
-These sparse vectors are used to cluster all examples.
+We first encode all examples into pre-trained embeddings using all-MiniLM-L6-v2, then use the trained MLP to derive BoT versions of these embeddings.
+These BoT vectors are used to cluster all examples.
 Dataset $K$ Total LLM No-LLM LLM ratio
-Bank77 77 13,083 5,780 7,303 44.2%
-Clinc150 150 22,500 7,650 14,850 29.6%
-Stackoverflow 20 20,000 6,400 13,600 32.0%
+Bank77 77 13,083 3,080 10,003 23.5%
+Clinc150 150 22,500 4,500 18,000 20%
 Table 16: Dataset Statistics.